### v1

# Scalable and Robust Multiband Modeling of AGN Light Curves in Rubin-LSST

Abstract
The Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST) will monitor tens of millions of active galactic nuclei (AGNs) for a period of 10 years with an average cadence of 3 days in six broad photometric bands. This unprecedented dataset will enable robust characterizations of AGN UV/optical variability across a wide range of AGN physical properties. However, existing tools for modeling AGN light curves are not yet capable of fully leveraging the volume, cadence, and multiband nature of LSST data. We present EzTaoX, a scalable light curve modeling tool designed to take advantage of LSST’s multiband observations to simultaneously characterize AGN UV/optical stochastic variability and measure interband time delays. EzTaoX achieves a speed increase of $\sim 10^{2}-10^{4}\times$ on CPUs over current tools with similar capabilities, while maintaining equal or better accuracy in recovering simulated variability properties. This performance gain enables continuum time-delay measurements for all AGNs discovered by LSST—both in the Wide Fast Deep survey and the Deep Drilling Fields—thereby opening new opportunities to probe AGN accretion-flow geometries. In addition, EzTaoX’s multiband capability allows robust characterization of AGN stochastic variability down to hourly timescales, facilitating the identification of accreting low-mass AGNs—such as those residing in dwarf galaxies—through their distinctive variability signatures.
software: pandas (McKinney, 2010), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), astropy (The Astropy Collaboration et al., 2013; The Astropy Collaboration et al., 2018; Astropy Collaboration et al., 2022), emcee (Foreman-Mackey et al., 2013), eztao (Yu & Richards, 2022), tinygp (Foreman-Mackey et al., 2024), numpyro (Bingham et al., 2019; Phan et al., 2019) eztaox (Yu & Burke, 2025)
1 Motivation
The UV/optical continuum luminosity of active galactic nuclei (AGNs) exhibit correlated stochastic variability on timescales ranging from hours to years (e.g., Vanden Berk et al., 2004; Sesar et al., 2007; Schmidt et al., 2012). This variability is believed to originate in the accretion disk, potentially driven in part by thermal reprocessing of X-ray photons emitted from the central corona (Krolik et al., 1991) and/or accretion rate fluctuations propagating through the disk (Lyubarskii, 1997). For a recent review of the observational properties of AGN variability in the UV/optical and X-ray bands, we refer the reader to Paolillo & Papadakis (2025).
A damped random walk (DRW) process is commonly adopted to model AGN UV/optical light curves (Kelly et al., 2009). The power spectral density (PSD) of a DRW exhibits a broken power-law (PL; $1/f^{\alpha}$) form, where the PL exponent $\alpha$ transitions from zero at low frequencies to two at high frequencies. The point at which this change in PL exponent occurs is known as the break frequency, which corresponds to a characteristic timescale denoted by $\tau_{\rm DRW}$. The asymptotic variability amplitude of the process is denoted by $\sigma_{\mathrm{DRW}}$. Although the DRW process is a phenomenological model and the precise PSD shape of intrinsic AGN UV/optical variability remains under debate (e.g., Mushotzky et al., 2011; Kasliwal et al., 2015; Smith et al., 2018; Arévalo et al., 2024; Beard et al., 2025; Petrecca et al., 2024; Yu et al., 2025), it continues to be a valuable and widely used tool for modeling AGN light curves because of its adequacy and computational efficiency.
Modeling AGN UV/optical light curves as a DRW has revealed compelling correlations between variability and AGN fundamental properties. Numerous previous investigations have shown that the best-fit $\sigma_{\mathrm{DRW}}$ correlates negatively with the Eddington ratio ($L/L_{\rm Edd}$) and bolometric luminosity ($L_{\rm bol}$) of AGNs (e.g., MacLeod et al., 2010; Sánchez-Sáez et al., 2018; Suberlak et al., 2021). Additionally, the characteristic timescale of the DRW model ($\tau_{\rm DRW}$) correlates positively with $M_{\rm{BH}}$ (e.g., Burke et al., 2021; Wang et al., 2023). These empirical relations not only provide valuable insights into the physical mechanisms driving AGN variability, but also form the basis for new techniques to estimate AGN fundamental properties using photometric data alone. For example, Burke et al. (2022) leveraged the $\tau_{\rm DRW}$$M_{\rm{BH}}$ correlation to search for low-mass AGNs powered by accreting massive black holes (MBH; $10^{5}\,$$M_{\odot}$$M_{\rm{BH}}$$10^{7}\,$$M_{\odot}$) at the centers of dwarf galaxies  (i.e., $M_{\star}\lesssim 10^{10}\,M_{\odot}$; Greene et al., 2020).
AGN UV/optical variability at longer wavelengths is also often observed to lag behind the variability at shorter wavelengths (Collier et al., 1999; Sergeev et al., 2005). Under the thermal reprocessing scenario, longer-wavelength photons originate from disk annuli located farther from the central supermassive black hole (SMBH) than those emitting shorter-wavelength photons (Shakura & Sunyaev, 1973). Consequently, the reprocessed emission at longer wavelengths exhibits longer light-crossing time delays relative to the central irradiating source (Cackett et al., 2007). This specific geometry—in which a compact source near the SMBH (e.g., an X-ray corona) irradiates the accretion disk—is commonly referred to as the lamp-post model. In this framework, the size of the accretion disk can be inferred by measuring the time lags between UV/optical light curves at different wavelengths—a technique known as continuum reverberation mapping (CRM), which probes the geometry of unresolved sources through time-delay measurements (see Cackett et al., 2021, and references therein).
Numerous CRM experiments have reported that AGN accretion disk sizes are systematically larger than those predicted by the standard thin-disk model (e.g., Fausnaugh et al., 2016; Jiang et al., 2017; Jha et al., 2022). Larger-than-expected accretion disk sizes have also been independently inferred from microlensing observations of gravitationally lensed quasars, where the disk half-light radius (i.e., the characteristic size) is inferred from the differential microlensing variability exhibited by the multiple lensed images (e.g., Pooley et al., 2007; Dai et al., 2010; Morgan et al., 2010; Jiménez-Vicente et al., 2014; Cornachione et al., 2020). It has been suggested that the measured lags may be contaminated by reprocessed UV/optical photons originating in the broad emission-line region (BLR), which lies farther out in the accretion flow (e.g., Korista & Goad, 2001; McHardy et al., 2018; Cackett et al., 2018; Lawther et al., 2018). This contamination could cause the observed interband continuum lags to appear longer than those expected from pure disk reprocessing.
In contrast, some studies have argued that accretion disk sizes inferred from CRM are consistent with the standard thin-disk model when factors such as temperature fluctuations within the disk or potential selection biases are properly accounted for (e.g., Dexter & Agol, 2011; Mudd et al., 2018; Homayouni et al., 2019; Yu et al., 2020b). Moreover, Kammoun et al. (2021) showed that the standard thin-disk model can reproduce the observed CRM lags when relativistic (special and general) and disk ionization effects are fully included. The accretion disk size discrepancy poses one of the significant challenges to the standard thin-disk model and to our understanding of AGN accretion physics more broadly. A large sample of AGNs with robustly measured CRM lags is required to investigate this issue further. To date, however, CRM lags have been successfully measured for only a few hundred AGNs (e.g., Collier et al., 1998; Peterson et al., 1998; Sergeev et al., 2005; Jiang et al., 2017; Homayouni et al., 2019; Mudd et al., 2018; Yu et al., 2020b; Guo et al., 2022).
[…]
The high-quality light curves provided by LSST will enable robust characterization of AGN UV/optical variability and interband continuum lags. Interband lags can be measured for about one million LSST AGNs (Yu et al., 2020a, Li et al. 2025 submitted), allowing detailed investigations of the accretion disk size problem as a function of AGN properties such as $L_{\rm Edd}$ and $M_{\rm BH}$, and offering new insights into the physical origin of continuum lags. Meanwhile, stochastic variability parameters—such as characteristic timescales and amplitudes—extracted from LSST light curves will enable reliable AGN classification (e.g., Kozłowski et al., 2010; Butler & Bloom, 2011; MacLeod et al., 2011; Peters et al., 2015; De Cicco et al., 2021; Sánchez-Sáez et al., 2021; Savić et al., 2023; Sánchez-Sáez et al., 2023). For example, variability-based methods are more effective than SED-based methods at identifying low-luminosity and low-mass AGNs (e.g., in dwarf galaxies), where host-galaxy starlight contamination is significant (e.g., Baldassare et al., 2017; Ward et al., 2022; Bernal et al., 2025). A complete census of low-luminosity is essential for understanding AGN evolution across cosmic time (Richards et al., 2006), while a comprehensive inventory of low-mass AGNs is critical for constraining the formation pathways of SMBHs (see Greene et al., 2020, and references therein).
However, existing AGN light curve analysis tools are not well-prepared to fully exploit the scale and richness of LSST observations. Current tools/methods used to characterize AGN UV/optical variability typically operate on single-band light curves and lack the ability to incorporate multiband information (e.g., Vaughan et al., 2003; Kelly et al., 2009, 2014; Kozłowski, 2016; Yu & Richards, 2022). This limitation prevents the effective utilization of LSST’s multiband coverage for improved characterization of AGN variability. Tools/methods commonly used to measure interband lags generally fall into two categories. The first includes techniques that estimate lags by computing the cross-correlation function (CCF) between light curves; the CCF reaches its maximum when one light curve is shifted relative to the other by an amount corresponding to the intrinsic lag (e.g., Gaskell & Peterson, 1987; Edelson & Krolik, 1988; Alexander, 1997). [^1] When applied to a single-band light curve and its shifted copies, the CCF method yields its autocorrelation function (ACF), which can be used to characterize AGN variability (e.g., Raiteri et al., 2021a, b). These methods are typically computationally efficient, but are limited to lag estimation and do not provide a full characterization of the underlying variability process. The second consists of methods that model each observed light curve as the convolution of an unknown driving continuum with a time-lag distribution, represented by a transfer function (e.g., Zu et al., 2011; Li et al., 2016; Starkey et al., 2016; Donnan et al., 2021). This is expressed as,
$y(t)=\int d\tau\,\Psi(\tau)\,y_{\rm drive}(t-\tau),$ (1)
where $y_{\rm drive}(t-\tau)$ denotes the flux of the underlying driving continuum at $t-\tau$, and $\Psi(\tau)$ represents the corresponding transfer function. JAVELIN is a representative tool in this category (Zu et al., 2013). It models the driving continuum as a DRW process and adopts a top-hat profile for the transfer function. While such methods provide a more comprehensive treatment by jointly modeling variability and interband lags, they are often computationally demanding.
Many recent investigations have turned to machine learning (ML) techniques to develop more scalable methods for analyzing AGN light curves (e.g., Fagin et al., 2024; Li et al., 2024). In these approaches, deep neural networks are trained on simulated light curves and are tasked with recovering the underlying variability characteristics and/or the physical parameters used to generate the simulations. However, the reliability of such methods hinges on the assumption that the models used to generate the training data accurately represent both the variability physics and the observational conditions (e.g., cadence, photometric uncertainty)—an assumption rarely satisfied in practice. For example, Li et al. (2024) show that the AGN physical parameters (e.g., $M_{\rm{BH}}$) predicted by their simulation-based inference model (Cranmer et al., 2020) can be biased by up to 0.5 dex when the test light curves are five times noisier than those used for training.
In this paper, we present EzTaoX—a scalable and flexible AGN light curve modeling tool that leverages LSST’s multiband observations to simultaneously infer the properties of the underlying stochastic variability and the lags between photometric bands (Yu & Burke, 2025). The name of EzTaoX combines three elements: “Ez", a homophone of easy; “Tao" referencing the Chinese philosophical and religious tradition of Taoism, commonly translated as “the Way"; and “X" denoting that the code is implemented in JAX—a high-performance numerical computing library developed by Google. EzTaoX achieves a speed increase of $\sim 10^{2}-10^{4}\times$ on CPUs compared to existing tools with similar capabilities (e.g., JAVELIN), making it feasible to perform both stochastic variability characterization and interband continuum lag measurements for all AGNs discovered by LSST. Furthermore, EzTaoX’s multiband capability enables robust constraints on stochastic variability timescales (e.g., $\tau_{\rm DRW}$) down to one hour with the LSST WFD cadence, facilitating the identification of low-mass AGNs—such as those residing in dwarf galaxies—which are expected to exhibit characteristic $\tau_{\rm DRW}$ ranging from hours to days (Burke et al., 2021; Wang et al., 2023).
The rest of the paper is organized as follows. Section 2 describes the methods and algorithms implemented in EzTaoX for characterizing AGN stochastic variability and interband continuum lags. Section 3 presents the performance of EzTaoX in terms of computational scalability and its robustness in recovering stochastic variability parameters and interband lags. In Section 4, we discuss the scientific opportunities enabled by EzTaoX, limitations of the current implementation, and potential directions for future development. Finally, we summarize our findings and conclude in Section 5.
2 Light Curve Modeling Method & Implementation
EzTaoX models AGN UV/optical stochastic variability as a Gaussian processes (GP), a method commonly adopted in AGN variability analysis (e.g., Press et al., 1992; Kelly et al., 2009; Kozłowski et al., 2010; Wilkins, 2019; Griffiths et al., 2021; Yu et al., 2022; Stone et al., 2022; Zhang et al., 2023; McLaughlin et al., 2024). Since the flux light curves of accreting compact objects typically follow a lognormal distribution (e.g., Uttley & McHardy, 2001; Gaskell, 2004; Giebels & Degrange, 2009; MacLeod et al., 2012; De Cicco et al., 2022), we emphasize that GP modeling is best applied to magnitude light curves, which follow a Gaussian distribution. [^2] Gúrpide & Middleton (2025) demonstrated that even when the light curves follow a lognormal distribution, the simulated GP parameters can be recovered without significant bias.
Given a single band AGN light curve, a GP treats the intrinsic brightness of the AGN at each point in time as a Gaussian random variable and captures the temporal correlations between observations through a kernel function $k(t_{i},t_{j})$, which defines the covariance between the Gaussian variables at any two timestamps $t_{i}$ and $t_{j}$ (Rasmussen & Williams, 2006; Aigrain & Foreman-Mackey, 2023).
The likelihood function of a GP given a set of observations $\bf{y}$ is,
$\ln\mathcal{L}$ $=-\frac{1}{2}(\mathbf{y}-\boldsymbol{\mu})^{\top}\mathbf{K}^{-1}(\mathbf{y}-\boldsymbol{\mu})$
$\quad\,-\frac{1}{2}\ln\det(\mathbf{K})-\frac{N}{2}\ln(2\pi)$
where $\boldsymbol{\mu}$ is the mean vector with its entries equal to the mean of $\bf{y}$, and $N$ is the total number of observations in $\bf{y}$. $\mathbf{K}$ is the GP covariance matrix, with each entry given by
$K_{ij}=k(|t_{j}-t_{i}|)+\sigma_{i}^{2}\,\delta_{ij},$ (3)
where $k$ is the aforementioned kernel function, and $t_{i}$ and $t_{j}$ are the timestamps of the $i^{\mathrm{th}}$ and the $j^{\mathrm{th}}$ observation, respectively. The parameter $\sigma_{i}$ denotes the measurement uncertainty of the $i^{\mathrm{th}}$ observation, and $\delta_{ij}$ is the Kronecker delta. The widely-adopted DRW model can expressed as a GP with the following kernel function:
$k_{\rm DRW}(|\Delta t|)=\sigma_{\rm DRW}^{2}\,e^{-|\Delta t|/\tau_{\rm DRW}},$ (4)
where $\Delta t$ gives the time difference between any two observations in a light curve.
EzTaoX performs multiband modeling and interband lag measurements by treating mean-subtracted light curves from different photometric bands as scaled and time-shifted realizations of a shared underlying GP. This approach is mathematically equivalent to adopting a Dirac delta function as the transfer function in Equation 1, with the delta function centered at a constant temporal offset. Under this framework, light curves from different photometric bands and different photometric surveys can be combined and modeled using a one-dimensional GP. This approach can in principle accommodate any number of photometric bands, although the number of free parameters grows with the number of distinct bands.
The corresponding kernel function for any two observations in such a merged light curve is,
$k(|t_{j}-t_{i}|)=S_{1}S_{2}\,k_{\rm latent}(|t_{j}-t_{i}-\tau_{\rm lag}|),$ (5)
where $k_{\rm latent}$ denotes the kernel of the underlying latent GP. $S_{1}$ and $S_{2}$ are the variability-amplitude scaling factors for the corresponding photometric bands, and $\tau_{\rm lag}$ represents the relative time lag between them. The variability amplitude in each band is obtained by multiplying the latent GP amplitude by the corresponding scaling factor. This class of kernel functions, as expressed in Equation 5, has seen broad application in astronomical time series analysis (e.g., Gordon et al., 2020; Villar et al., 2021). To evaluate the likelihood of the multiband model for a given light curve, we shift the light curve in different bands by their trial $\tau_{\rm lag}$ relative to the reference band, and compute Equation 2 using a kernel that assumes no intrinsic lag between bands (i.e., setting $\tau_{\rm lag}=0$ in Equation 5). In this way, the GP parameters and the interband lags are fitted jointly.
EzTaoX is built on top of tinygp [^3] https://github.com/dfm/tinygp—a GP modeling framework developed using JAX. [^4] https://github.com/jax-ml/jax JAX is a high-performance numerical computing library that combines automatic differentiation, just-in-time (JIT) compilation, and hardware acceleration to empower scalable machine learning and scientific computing. tinygp leverages these features of JAX and re-implements the novel celerite algorithm introduced by Foreman-Mackey et al. (2017) in JAX. When a kernel function can be expressed as a mixture of exponentials, the celerite algorithm can evaluate the GP likelihood function with $\mathcal{O}(N)$ computational scaling, a significant efficiency enhancement over the standard $\mathcal{O}(N^{3})$ scaling associated with direct likelihood evaluation (Foreman-Mackey et al., 2017). EzTaoX takes advantage of the JAX-based implementation of the celerite algorithm to provide a scalable solution for multiband light curve modeling of AGNs.
PSD shapes exhibited by the light curves of accreting black holes can be effectively modeled with celerite GP kernels. For example, the continuous-time autoregressive moving-average (CARMA) process—commonly used as a more flexible alternative to the DRW process—has an equivalent celerite implementation (Roux, 2002; Kelly et al., 2014; Kasliwal et al., 2017; Moreno et al., 2019; Yu et al., 2022). A CARMA($p$, $q$) process combines an autoregressive process of order $p$ with a moving-average process of order $q$, and its PSD can be expressed as a weighted sum of modified Lorentzians. The DRW process is a special case of a CARMA process with $p=1$ and $q=0$. For a detailed description of CARMA modeling, we refer the reader to Kelly et al. (2014). More recently, Lefkir et al. (2025) introduced an algorithm for approximating arbitrary broken PL PSDs using a sum of multiple celerite kernels, thereby enhancing the interpretability and flexibility of GP-based modeling for AGN light curves. Moreover, celerite-based kernels can be freely combined through addition and multiplication to model more complex covariance structures, while retaining the $\mathcal{O}(N)$ computational scaling. All such kernels can be used for simultaneous multiband variability characterization and interband lag measurements.
Table 1: Input Parameters for Simulated LSST WFD AGN Light Curves
Light Curve Set $g$-band $\tau_{\rm DRW}$ $g$-band $\sigma_{\mathrm{DRW}}$ Interband Lag Band Duration
[day] [mag] [day] [year]
One 100 0.112 Uniform(0.01, 10) $g$, $r$ 3, 10
[…]
Three 100 LogUniform(0.02, 0.2) 1 $u$, $g$, $r$ 3
3 EzTaoX Performance
In this section, we evaluate the robustness and scalability of EzTaoX and compare it to the widely adopted tool JAVELIN. Since JAVELIN supports only the DRW kernel for modeling the driving continuum variability during lag measurement, all performance experiments conducted in this work assume a DRW model for the underlying AGN continuum stochastic variability. Section 3.1 demonstrates EzTaoX’s robustness in recovering input interband lags and DRW parameters from simulated LSST WFD light curves. In Section 3.2 and Section 3.3, we apply EzTaoX to continuum reverberation mapping light curves of real AGNs and compare the resulting lag measurement with values reported in the literature. Finally, Section 3.4 presents computational benchmarking results for EzTaoX.
3.1 Robustness: A Test Using Simulated LSST WFD Light Curves
In this section, we simulate realistic LSST WFD light curves, and assess how well EzTaoX recovers both the simulated interband lags and stochastic variability parameters. Section 3.1.1 describes the light curve simulation process and the sets of input parameters (see Table 1). In Section 3.1.2, we show that EzTaoX is capable of retrieving interband lags with accuracy comparable to that of JAVELIN. Section 3.1.3 and Section 3.1.4 demonstrates the robustness of EzTaoX in recovering input stochastic variability parameters (i.e. $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$). We emphasize that the goal of the experiments in this section is to demonstrate the overall capability of EzTaoX, rather than to perform a full parameter-space study or to investigate its systematics. Nonetheless, the code for generating and fitting simulated LSST light curves will be made available upon request, allowing interested reader to further explore the robustness of EzTaoX under different combinations of input parameters.
3.1.1 Light Curve Simulation
We adopt the DRW model to simulate AGN light curves in individual LSST bands. We start by simulating DRW light curves in each LSST band with a 30-second cadence; this 30-second cadence is chosen to match the 30-second exposure time of each LSST single visit (Ivezić et al., 2019). The input DRW parameters across different bands are correlated according to the wavelength-dependent power-law relations presented in MacLeod et al. (2010). More specifically,
$\tau_{\rm DRW}\propto\lambda^{0.17},\quad\sigma_{\rm DRW}\propto\lambda^{-0.479},$ (6)
where $\lambda$ is the effective wavelength of the LSST filters (Ivezić et al., 2019). We then simulate continuum lags by offsetting light curves in redder band in time with respect to light curves in bluer bands. Finally, we down-sample the simulated 30-second cadence light curve to the observing cadence of the WFD survey.
The observational noise is simulated by adding random values drawn from Gaussian distributions with mean zero and variance $\sigma^{2}_{\mathrm{LSST}}$. The parameter $\sigma_{\mathrm{LSST}}$ is the expected photometric error of LSST, and it can be approximated using
$\sigma^{2}_{\mathrm{LSST}}=\sigma^{2}_{\mathrm{sys}}+\sigma^{2}_{\mathrm{rand}},$ (7)
where $\sigma_{\mathrm{sys}}$ is the systematic photometric error and $\sigma_{\mathrm{rand}}$ is the random photometric error. LSST aims to achieve a $\sigma_{\mathrm{sys}}<0.005$ mag (Ivezić et al., 2019); we set $\sigma_{\mathrm{sys}}$ to $0.005$ mag in our simulation. The random photometric error ($\sigma_{\mathrm{rand}}$) is expected to scale as
$\sigma^{2}_{\mathrm{rand}}=(0.04-\gamma)x+\gamma x^{2}\,(\mathrm{mag}^{2}),\quad x=10^{0.4(m-m_{5})},$ (8)
[…]
The observing cadence and the expected $m_{5}$ at any location on the sky are predicted by LSST operation simulations (OpSims) using MAF (Marshall et al., 2017). In this work, we adopt the operation simulation version baseline_v4.0_10yrs. All our simulated AGNs also assume a fiducial $g$-band mean magnitude of 21 ($\sim 1\%$ photometric accuracy) and mean uugg and ggrr colors of 0.1 mag (Richards et al., 2001; Temple et al., 2021). An example light curve from the first set is shown in Figure 1.
Figure 1: Left: Example LSST WFD light curves simulated using the DRW model. The $g$- and $r$-band light curves have fiducial mean magnitudes of 21 and 20.9, respectively. Right: The true (dashed curves) and recovered (colored solid curves) PSDs. The recovered PSDs are generated using random draws of the MCMC posterior distribution. The orange histogram in the bottom right panel shows the posterior distribution of the recovered ggrr lag, with the dash-dotted vertical line indicating the true simulated lag of 2 days.
We generate three sets of multiband LSST light curves using different input distributions for the interband lags and DRW parameters. The distributions of input parameters are also summarized in Table 1. Each set consists of 975 simulated light curves with their cadence and $m_{5}$ retrieved from 975 unique locations uniformly sampled in the WFD footprint. To construct this sample, we first generate a uniform grid with 5 spacing in both RA and Dec, spanning 0 << RA << 360 and $75^{\circ}$ << Dec << $15^{\circ}$. We then select grid points that receive between 600 and 1000 visits in all six photometric bands. The final simulated full 10-year light curves have median numbers of 53, 65, and 171 epochs in the $u$, $g$, and $r$ bands, respectively. When only the first three years of data are considered, these numbers drop to 13, 19, and 50 epochs.
The first set aims to evaluate EzTaoX’s ability to recover interband lags (see Section 3.1.2), and the light curves consist of simulated LSST observations in $g$ and $r$ bands. The ggrr lag for each multiband light curve is uniformly sampled between 0.01 day and 10 days, while the input $g$-band $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ are fixed at fiducial values of 100 days and 0.112 mag, respectively. These choices of $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ are representative of those measured from typical quasar (i.e., high luminosity AGN) UV/optical light curves (e.g., Vanden Berk et al., 2004; MacLeod et al., 2010).
The second set tests EzTaoX’s ability to recover the DRW characteristic timescale $\tau_{\rm DRW}$ (see Section 3.1.3) , and the light curves consist of simulated LSST observations in $u$, $g$, and $r$ bands. The input $g$-band $\tau_{\rm DRW}$ is randomly sampled from a log-uniform distribution between 0.01 day and 500 days, while the $g$-band $\sigma_{\mathrm{DRW}}$ is set to a fiducial value of 0.112 mag. The $u$-$g$ and ggrr lags are set to 1% of the $g$-band $\tau_{\rm DRW}$.
The third set evaluates EzTaoX’s ability to recover the DRW asymptotic variability amplitude $\sigma_{\mathrm{DRW}}$ (see Section 3.1.4), and the light curves consist of simulated LSST observations in $u$, $g$, and $r$ bands. The input $g$-band $\sigma_{\mathrm{DRW}}$ is randomly sampled from a log-uniform distribution between 0.02 mag and 0.2 mag, while the $g$-band $\tau_{\rm DRW}$ is set to a fiducial value of 100 days. The $u$-$g$ and ggrr lags are set to 1 day.
3.1.2 Recovery of Interband Lag
Figure 2: Comparison of ggrr lags recovered using JAVELIN (left) and EzTaoX (right) from simulated LSST WFD light curves. Both JAVELIN and EzTaoX yield comparable results in terms of lag measurement accuracy. Top: The difference between the recovered and true lags ($\Delta\>\rm lag$) as a function of the true lag, based on fits obtained using only the first 3 years of WFD observations. The dispersion of $\Delta\,\rm lag$, computed using the median absolute deviation (MAD), is indicated in the top-left corner of each panel. Bottom: Same as the top panels, but using the full 10 years of WFD observations for inference.
Table 2: MCMC Prior for EzTaoX Parameters
Parameter Prior Distribution
$\tau_{\rm DRW}$ [day] LogUniform($10^{-4}$, $10^{4}$)
$\sigma_{\mathrm{DRW}}$ [mag] LogUniform($0.01$, 22)
$S_{\rm band}$ LogUniform(0.1{0.1}, 10{10})
$\tau_{\rm lag}$ [day] Uniform(15-15, $25$)
$\mu_{\rm band}$ [mag] Normal(0, $0.1$)
Note. — The normal distribution follow the notation Normal($\mu$, $\sigma$), where $\mu$ is the mean, and $\sigma$ is the width.
We run EzTaoX on the first set of simulated light curves to extract the lag between the $g$- and $r$-band ($\tau_{\rm gr}$), after subtracting the mean from each single-band light curve. The EzTaoX light curve model includes seven parameters: $\tau_{\rm DRW}$, $\sigma_{\mathrm{DRW}}$, $S_{\rm g}$, $S_{\rm r}$, $\mu_{\rm g}$, $\mu_{\rm r}$, and $\tau_{\rm gr}$. Here, $S_{\rm g}$ and $S_{\rm r}$ are the GP kernel scaling factors for $g$- and $r$-band observations (see Equation 5), while $\mu_{\rm g}$ and $\mu_{\rm r}$ are the mean magnitudes of the $g$- and $r$-band light curves, respectively. As part of the EzTaoX fitting process, we first compute the maximum likelihood estimates (MLEs) of the model parameters, and use the them to initialize the Markov Chain Monte Carlo (MCMC) sampler for posterior sampling. The adopted MCMC priors are listed in Table 2 and are fixed throughout the analysis unless stated otherwise. The prior ranges are chosen according to the light curve properties. For example, the lower limit for $\tau_{\rm DRW}$ is set to $\sim$10 times smaller than the minimum time separation between any two observations, while the upper limit is $\sim$10 times larger than the light curve length. MCMC sampling is performed using the gradient-based No-U-Turn Sampler (NUTS; Hoffman & Gelman, 2011), with a single chain run for $10^{4}$ steps following $10^{3}$ warm-up steps. The posterior median of $\tau_{\rm DRW}$ is taken as the point estimate.
We also apply JAVELIN to the first set of simulated light curves to measure $\tau_{\rm gr}$. The JAVELIN light curve model includes five parameters: $\tau_{\rm DRW}$, $\sigma_{\mathrm{DRW}}$, $s_{\rm gr}$, $w_{\rm gr}$, and $\tau_{\rm gr}$. Here, $s_{\rm gr}$ and $w_{\rm gr}$ describe the scale and width, respectively, of the top-hat transfer function for the ggrr lag. JAVELIN employs a two-step fitting procedure: it first generates a distribution of $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ by fitting the light curve of a chosen band via MCMC sampling, and then uses that distribution as a prior for the DRW parameters while simultaneously fitting both the DRW parameters and the interband lags from the light curves in all bands. In the first step, the light curve in the bluest band is typically used, as JAVELIN assumes that redder-band light curves are convolved versions of it; here we use the $g$-band light curve. For MCMC sampling, JAVELIN employs the affine-invariant ensemble sampler (Goodman & Weare, 2010) as implemented in emcee (Foreman-Mackey et al., 2013). In the first step we follow the default configuration, while in the second step we run the sampler with 100 walkers, 500 burn-in steps, and 3000 sampling steps. We also set the prior bound for $\tau_{\rm DRW}$ to $10^{-4}$, $10^{4}$] days and for $\tau_{\rm gr}$ to [15-15, $25$] days. The posterior median of $\tau_{\rm gr}$ is reported as the final point estimate.
We clean the sample of best-fit EzTaoX and JAVELIN lags by removing those associated with non-convergent MCMC chains. Specifically, we discard fits in which the MCMC chain for $\tau_{\rm gr}$ yields fewer than 400 effective samples ($n_{\rm eff}$), following the recommendation of Vehtari et al. (2021) for selecting converged chains. Approximately 20% of the 3-year and 7% of the 10-year EzTaoX fits are removed by this criterion, while $\sim$20% of the JAVELIN fits are removed from both the 3-year and 10-year data. Examined as a function of the input $\tau_{\rm gr}$, the fraction of discarded JAVELIN fits increases with increasing true $\tau_{\rm gr}$; in contrast, the fraction of discarded EzTaoX fits remains relatively constant across the range of input $\tau_{\rm gr}$. We discuss this effect in more detail in Section 4.3.
[…]
Based on the cleaned samples, EzTaoX and JAVELIN yield comparable results for lag measurement accuracy. Figure 2 presents the difference between the recovered lag and the true simulated lag ($\Delta\rm lag$) as a function of the true lag. The left column shows results from JAVELIN, while the right column displays results from EzTaoX. The top row corresponds to fits obtained using only the first 3 years of simulated LSST light curves, and the bottom row shows results based on the full 10-year simulated dataset. The dispersion in $\Delta\rm lag$ is $\lesssim\>$2 days when using 3 years of data and improves to $\lesssim\>$0.7 day with the full 10-year dataset. These dispersions are consistent with the mean uncertainties derived from the MCMC posterior distributions.
3.1.3 Recovery of Stochastic Variability Timescale
We follow the same strategy described in the previous section to fit the second set of simulated light curves. We apply EzTaoX to multiband light curves containing observations in $u$, $g$, and $r$ bands, as well as to light curves with observations in only one of the three bands. We use the same set of MCMC configurations as those employed in Section 3.1.2. When fitting the multiband light curves, we observed that MCMC chains take significantly longer to converge when the true interband lag is shorter than a few hours. This behavior may be driven by the minimum interband cadence (i.e., uugg, ggrr, etc.) of $\sim 40$ minutes, which hampers reliable lag measurements on hourly timescales and below. To mitigate this issue, and without loss of generality, we fix the interband lags to zero when the MLE of $\tau_{\rm DRW}$ fall below 10 days—corresponding to a simulated lag of $\sim 2.5$ hours. We exclude fits that either exhibit poorly-converged MCMC chains ($n_{\rm eff}<400$) or yield a best-fit $\tau_{\rm DRW}$ more than an order of magnitude smaller than the minimum cadence. About 5% of EzTaoX fits are removed by these two criteria.
[…]
Figure 4: Comparison of recovered and input $\tau_{\rm DRW}$ from simulated LSST WFD light curves without hourly revisits. When the input $\tau_{\rm DRW}$ is less than a few days, estimates recovered using only the $u$- or $g$-band observations show systematic bias. In contrast, combining both $u$- and $g$-band observations yields unbiased $\tau_{\rm DRW}$ estimates down to hourly timescales.
To better demonstrate the advantages of multiband fitting, we refit the second set of simulated light curves after removing revisits occurring within 1 hr in any given band. This removal mimics a scenario in which intra-night observations must be stacked to detect faint variable sources against host-galaxy starlight—such as low-luminosity and low-mass AGNs, which are expected to exhibit shorter characteristic variability timescales than typical quasars (e.g., Burke et al., 2021). Figure 4 displays the recovered $\tau_{\rm DRW}$ as a function of the input value. When using only $u$- or $g$-band observations, the recovered $\tau_{\rm DRW}$ is systematically biased for input values below a few days. In contrast, combining $u$- and $g$-band data allows $\tau_{\rm DRW}$ to be recovered without bias down to hourly timescales. Furthermore, MCMC chains fail to converge (i.e., $n_{\rm eff}<400$) in more than 40% of the single-band fits, whereas this fraction decreases to only 7% for the multiband fits, underscoring the improved reliability of multiband fitting. The multiband capability of EzTaoX thus enables robust selection of low-luminosity and low-mass AGNs based on variability, particularly those with intrinsically faint emission and stochastic variability characterized by $\tau_{\rm DRW}$ values of just hours to days.
We also run JAVELIN on the second set of light curves with observations in all three bands. To improve convergence, we increase the number of MCMC sampling steps from 3000 to $10^{4}$. In the first fitting step, we use the $r$-band light curve to generate the prior distribution of $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$, since the $r$-band provides the densest temporal sampling and lowest photometric uncertainty. As shown in Appendix B, the final DRW parameter distributions recovered by JAVELIN depend on the photometric band used in the first step, with the $r$-band yielding the most reliable results. As with EzTaoX, we find that the MCMC chains exhibit poor convergence when the true interband lag is shorter than a few hours. To mitigate this issue, and without loss of generality, we fix the interband lags to zero when the MLE of $\tau_{\rm DRW}$ falls below 10 days. Finally, we apply the same sample-cleaning cuts used for EzTaoX, which remove about 5% of the JAVELIN fits.
[…]
3.1.4 Recovery of Stochastic Variability Amplitude
We fit the third set of light curves with observations in all three bands using both EzTaoX and JAVELIN, adopting the same priors and MCMC setup as in Section 3.1.3. We exclude fits that either exhibit poorly-converged MCMC chains ($n_{\rm eff}<400$) or yield a best-fit value of $\sigma_{\rm DRW}<0.003$ mag, an order of magnitude below the simulated photometric uncertainty (e.g., $\sigma_{{\rm LSST:}u}\approx 0.03$ mag). This removes $\sim$5% of the EzTaoX and JAVELIN fits.
Figure 5: Comparison of recovered and input $\sigma_{\mathrm{DRW}}$ from multiband light curves using $u$-, $g$-, and $r$-band observations with EzTaoX and JAVELIN. The hatched yellow region and the purple region represent the central 68% interval of EzTaoX- and JAVELIN-inferred $\sigma_{\mathrm{DRW}}$, respectively. EzTaoX yields unbiased estimates of $\sigma_{\mathrm{DRW}}$ across the full simulated range, from levels comparable to the photometric uncertainty up to $\sim\,$0.2 mag. In contrast, JAVELIN fails to recover $\sigma_{\mathrm{DRW}}$ when it approaches a few times the $u$-band photometric uncertainty, as marked by the vertical dash-dotted line.
EzTaoX is able to recover $\sigma_{\mathrm{DRW}}$ without bias down to levels comparable to the photometric uncertainty of the simulated observations. Figure 5 shows the recovered versus input $\sigma_{\mathrm{DRW}}$. The hatched yellow region and the purple region denote the central 68% intervals of $\sigma_{\mathrm{DRW}}$ recovered by EzTaoX and JAVELIN, respectively. Across the full range of simulated $\sigma_{\mathrm{DRW}}$, EzTaoX provides unbiased estimates. In contrast, JAVELIN systematically underestimates $\sigma_{\mathrm{DRW}}$, with values falling below the one-to-one relation (dashed diagonal). We suspect the bias likely arise from the underestimation of $\tau_{\rm DRW}$ seen in the middle panel of Figure 3. Because $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ are correlated, JAVELIN yields biased estimates of both parameters when the true $\tau_{\rm DRW}$ exceeds $\sim$30 days in a 3-year light curve. Finally, to demonstrate the robustness of EzTaoX in this regime, we simulate and fit an additional set of light curves with $\tau_{\rm DRW}$ =200=200 days; the results are presented in Appendix A.
Figure 6: Comparison of interband continuum lags for NGC 5548 measured with EzTaoX to those reported in the literature. The light curves of NGC 5548 are obtained by the AGN STORM project. In the legend, F16 refers to the lags reported by Fausnaugh et al. (2016), and E15 refers to the lags reported by Edelson et al. (2015). Interband continuum lags measured with EzTaoX are broadly consistent with those obtained using JAVELIN and the ICCF method, highlighting the effectiveness of EzTaoX despite its adoption of a simplified transfer function (i.e., a Dirac delta function).
Figure 7: Comparison of interband continuum lags for a sample of type 1 AGNs measured with EzTaoX to those reported in the literature (Guo et al., 2022). The light curves are obtained from ZTF. Interband continuum lags measured with EzTaoX are broadly consistent with those obtained using JAVELIN (left) and the ICCF method (right)
3.2 A Test Using Continuum Reverberation Mapping Light Curves of NGC 5548
The AGN Space Telescope and Optical Reverberation Mapping (STORM) project is an intensive multiwavelength monitoring campaign targeting NGC 5548, a well-studied Seyfert galaxy (De Rosa et al., 2015; Edelson et al., 2015; Fausnaugh et al., 2016). This campaign combines observations from the Hubble Space Telescope (HST), the Neil Gehrels Swift Observatory (Swift hereafter; Gehrels et al., 2004), and a network of ground-based telescopes, to provide simultaneous daily light curves of NGC 5548 across the X-ray, UV, and optical bands. For further details on the observational setup and data, we refer interested readers to De Rosa et al. (2015), Edelson et al. (2015), and Fausnaugh et al. (2016).
Applying EzTaoX to the UV/optical AGN STORM light curves of NGC 5548 reveals that the lags inferred by EzTaoX are broadly consistent with literature values derived using JAVELIN and ICCF. We follow the same strategy described in Section 3.1.2 to measure interband lags relative to the continuum light curve at 1367Å. In contrast to the approaches taken by Edelson et al. (2015), and Fausnaugh et al. (2016), we do not de-trend the light curves prior to lag estimation. We also increased the prior range of $S_{\rm band}$ to Uniform($^{-5}, 1010$) because AGN STORM light curve are provided in flux, which exhibits large amplitude variation across bands.
Figure 6 compares the lags measured with EzTaoX to those derived with JAVELIN and ICCF. F16 and E15 in the legend refer to the lags reported by Fausnaugh et al. (2016) and Edelson et al. (2015), respectively. The blue points represent the continuum lags measured at 1158Å, 1479Å, and 1746Å  using HST light curves. The green points correspond to the continuum lags from the Swift UVW2, UVM2, UVW1, U, and B bands. The yellow points represent continuum lags from Johnson/Cousins BVRI bands and Sloan Digital Sky Survey (SDSS; York et al., 2000) ugrizugriz bands based on ground-based observations. The purple squares in the right panel of Figure 6 shows the Swift lags reported by Edelson et al. (2015). We note that the Swift ICCF lags from Fausnaugh et al. (2016) are systematically larger than those reported by Edelson et al. (2015), likely due to the different de-trending approaches they adopted. Overall, interband lags measured using EzTaoX are broadly consistent with those obtained using JAVELIN and ICCF.
[…]
3.3 A Test Using Zwicky Transient Facility AGN Light Curves
[…]
We apply EzTaoX to the $g$- and $r$-band ZTF light curves of the core-sample AGNs from Guo et al. (2022), and find that the lags measured with EzTaoX are broadly consistent with those obtained using JAVELIN and ICCF. Figure 7 compares the interband lags measured with EzTaoX to those from JAVELIN (left panel) and ICCF (right panel). We retrieve the ZTF light curves for these sources from the IPAC Infrared Science Archive (ZTF Team, 2025) and truncate them to match the DR7 temporal baseline (i.e., MJD < 59400). When multiple light curves are available for the same object (each obtained from a different CCD), we select the one with the largest number of epochs. We follow the same strategy described in Section 3.1.2 to measure the lags between the $g$ and $r$ bands. We also fit additional jitter terms—one for each band—added to the diagonal elements of the covariance matrix to account for any residual photometric uncertainties not captured by the photometric pipeline (Burke et al., 2021).
3.4 Scalability: Computation Time
[…]
4 Discussion
4.1 Unbiased Point Estimator of $\tau_{\rm DRW}$
Unbiased recovery of $\tau_{\rm DRW}$ depends on the light-curve cadence, baseline (i.e., the total duration of observations), and adopted priors. As shown by Kozłowski (2017) and later studies (e.g., Sánchez-Sáez et al., 2018; Burke et al., 2021; Suberlak et al., 2021; Hu et al., 2024), $\tau_{\rm DRW}$ is systematically underestimated when the light-curve baseline is shorter than $\sim$10 times the true $\tau_{\rm DRW}$. Hu et al. (2024) conducted extensive simulations demonstrating that the assumed prior on $\tau_{\rm DRW}$ also impact its recovery. In addition, they reported that $\tau_{\rm DRW}$ is systematically overestimated when the true $\tau_{\rm DRW}$ is less than five times the mean cadence, since the temporal sampling then probes only the white-noise regime (i.e., a PL exponent of zero) of the DRW PSD.
Figure 9: Comparison of three point estimators for $\tau_{\rm DRW}$: MLE, MCMC posterior median, and MCMC posterior mean. The $x$-axis is the recovered $\tau_{\rm DRW}$-to-baseline ratio ($\rho_{\rm out}$), and the $y$-axis is the difference between the recovered and true values of log($\tau_{\rm DRW}$). In each panel, the dash-dotted line marks the median of the $\Delta\,$log($\tau_{\rm DRW}$) distribution, and the shaded region shows its central 68% interval. The MCMC posterior median is the most reliable, yielding systematically unbiased estimates up to one tenth of the light-curve baseline (vertical dotted line, left) and introducing only a small systematic bias of $\sim 0.1$ dex at one third of the baseline (vertical dotted line, right).
Following previous investigations, we perform simulations to assess how accurately EzTaoX recovers $\tau_{\rm DRW}$ as a function of $\tau_{\rm DRW}$-to-baseline ratio. We simulated 5000 single-band light curves with a uniform 2-day cadence and a baseline of $10^{4}$ days. Each light curve adopts a fiducial $\sigma_{\mathrm{DRW}}$ of 0.112 mag and a mean photometric uncertainty of 0.01 mag. The input $\tau_{\rm DRW}$ ranges from 10 days to $10^{4}$ days, corresponding to an input $\tau_{\rm DRW}$-to-baseline ratios ($\rho_{\rm in}$) spanning from 0.001 to 1. We fit the simulated single-band light curves using EzTaoX to infer $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$, following the strategy outlined in Section 3.1.2. To avoid potential prior-driven biases, we extend the upper limit of the MCMC prior on $\tau_{\rm DRW}$ to $10^{5}$ days—an order of magnitude longer than the light-curve baseline.
Figure 10: Left: Example LSST WFD light curves simulated using a CARMA(2,1) model. The $g$- and $r$-band light curves have fiducial mean magnitudes of 21 and 20.9, respectively. Right: The true (dashed curves) and recovered (colored solid curves) PSDs. The recovered PSDs are generated using random draws of the MCMC posterior distribution. As a comparison, the blue dashed curves display the reference DRW PSDs. The orange histogram in the bottom right panel shows the posterior distribution of the recovered ggrr lag, with the dash-dotted vertical line indicating the true simulated lag of 2 days.
Based on our simulations, EzTaoX recovers $\tau_{\rm DRW}$ without systematic bias up to one tenth of the light-curve baseline. Figure 9 displays the distribution of $\Delta\,$log($\tau_{\rm DRW}$)—the difference between the recovered and input log($\tau_{\rm DRW}$)—as a function of the recovered $\tau_{\rm DRW}$-to-baseline ratio ($\rho_{\rm out}$). Plotting $\rho_{\rm out}$ rather than $\rho_{\rm in}$ is more informative, since the true $\tau_{\rm DRW}$ is unknown when fitting DRW models to real AGN light curves. In each panel, the dash-dotted line indicates the median of the $\Delta\,$log($\tau_{\rm DRW}$) distribution, and the shaded region shows its central 68% interval. The top, middle, and bottom panels correspond to three common point estimators: MLE, MCMC posterior median, and MCMC posterior mean, respectively. Among these, the MCMC posterior median is the most robust, remaining systematically unbiased as long as the the recovered $\tau_{\rm DRW}$ is less than one-tenth of the baseline (dash-dotted lines in the middle panel of Figure 9). It introduces only a small systematic bias ($\sim 0.1$ dex) when the recovered $\tau_{\rm DRW}$ is about one-third of the baseline.
4.2 Deviation from DRW
In this work, we assume that AGN stochastic variability follows a DRW process; however, the modeling framework introduced here is sufficiently flexible to accommodate alternative GP kernels for characterizing the underlying variability. As an illustration, Figure 10 presents $g$- and $r$-band light curves simulated using a CARMA(2,1) process, sampled according to the LSST WFD cadence (left), along with the corresponding PSDs recovered using EzTaoX (right). The CARMA(2,1) process—also known as the noise-driven damped harmonic oscillator (DHO) model—has been shown to provide superior fits to observed AGN UV/optical light curves compared to the DRW model (e.g., Kasliwal et al., 2017; Yu et al., 2025). For reference, the DRW PSD is shown in blue in the right panel. Similar to Figure 1, the orange histogram in the lower right corner displays the posterior distribution of the ggrr lag. The specific CARMA(2,1) process simulated is governed by the following kernel:
$k(t_{i},t_{j})=a_{1}e^{-c_{1}\,|t_{j}-t_{i}|}+a_{2}e^{-c_{2}\,|t_{j}-t_{i}|},$ (9)
which is equivalent to the sum of two DRW kernels.
4.3 Beyond the Dirac Delta Transfer Function
EzTaoX models interband lags in multiband light curves using a Dirac delta transfer function, a simplified approach compared to methods that adopt more flexible transfer functions with finite width and skewed shapes (e.g., Zu et al., 2011; Starkey et al., 2016; Li et al., 2016; Donnan et al., 2021). While more sophisticated transfer functions can provide additional insights into the geometry and physical properties of AGN accretion disks (e.g., Cackett et al., 2007; Starkey et al., 2016; Fagin et al., 2024), we argue that the cadence of LSST WFD light curves is likely insufficient to resolve the detailed shape of the transfer functions. Thus, the simplified Dirac delta transfer function adopted by EzTaoX is likely adequate for lag estimation from LSST WFD light curves. Furthermore, previous investigations and the results of Section 3.2 show that even with high-cadence continuum reverberation mapping data, the use of more flexible transfer functions does not necessarily result in significantly different mean lag estimates compared to simpler methods (e.g., Fausnaugh et al., 2018; Donnan et al., 2023). An exception perhaps arises when the intrinsic transfer function is notably skewed, in which case simplified methods may underestimate the mean lag (Yu et al., 2020a; Chan et al., 2020).
As an extension to the comparison in Section 3.1.2, we test EzTaoX and JAVELIN on new sets of 10-year LSST WFD light curves simulated with a Gaussian transfer function. As before, the $g$-band light curves adopt an input $\tau_{\rm DRW}$ = 100 days and $\sigma_{\mathrm{DRW}}$ = 0.112 mag. The $r$-band light curves are generated by convolving the $g$-band curves with a Gaussian transfer function. We simulate three sets of light curves with fixed Gaussian widths of 0.5, 2, and 5 days. Following the same procedure as in Section 3.1.2, we extract ggrr lags from the simulated light curves using both EzTaoX and JAVELIN; the results are shown in Figure 11. We find that EzTaoX yields lag measurements comparable to those of JAVELIN even when the transfer function width is as large as 5 days. Moreover, the $n_{\rm eff}$-based sample cleaning procedure tends to removes a higher fraction of EzTaoX fits with true lags much smaller than the transfer function width, while preferentially discarding JAVELIN fits with true lags much larger than the width. This behavior explains the trend noted in Section 3.1.2, where the fraction of discarded JAVELIN fits increases with true lag when the light curves are simulated with a Dirac delta transfer function.
Figure 11: Comparisons of ggrr lags extracted by EzTaoX and JAVELIN from 10-year LSST WFD light curves simulated with Gaussian transfer functions of width 0.5, 2, and 5 days (top to bottom). The shaded regions shows the central 68% interval of the $\Delta\,$lag distribution. EzTaoX yields lag measurements comparable to those of JAVELIN, even when the light curves are simulated with a transfer function width as wide as 5 days.
4.4 LSST Variability Science Enabled by EzTaoX
EzTaoX enables reliable and simultaneous characterization of variability and interband lags for the tens of millions of AGNs to be monitored by LSST. Existing non-ML methods/tools capable of jointly characterizing stochastic variability and measuring interband lags are not scalable to LSST data volumes (e.g., Zu et al., 2011; Starkey et al., 2016; Donnan et al., 2021). While ML approaches are scalable, their reliability critically depends on the assumption that the training data accurately capture the true nature of AGN variability—an assumption that is rarely satisfied in practice. In comparison to traditional non-ML tools, EzTaoX achieves a speed increase of $100\times$ for LSST WFD light curves, and $10^{4}\times$ for LSST DDF light curves, without compromising —and in some cases improving—the accuracy of the results. Furthermore, EzTaoX offers the flexibility to incorporate more expressive GP kernels, allowing for improved modeling of intrinsic continuum variability. This combination of scalability and flexibility enables robust characterization of variability and interband lags for an unprecedented number of AGNs. A large, statistically significant sample of AGNs with well-characterized continuum variability and interband lags will facilitate detailed investigations into the fundamental drivers of AGN variability and the physical origins of continuum lags.
EzTaoX enables robust variability selection of actively accreting MBHs. Accreting MBHs in dwarf galaxies are difficult to identify due to their low luminosity relative to their host galaxy starlight (Greene et al., 2020). Nevertheless, their characteristic stochastic variability provides a powerful selection tool in the time domain (e.g., Baldassare et al., 2018; Ward et al., 2022; Burke et al., 2022; Bernal et al., 2025). By modeling dwarf galaxy image-differencing light curves as a DRW process, one can identify MBH candidates, which are expected to exhibit characteristic $\tau_{\rm DRW}$ values ranging from hours to days (Burke et al., 2021; Wang et al., 2023). Given the cadence of LSST WFD light curves, the multiband modeling capability of EzTaoX enables precise measurement of $\tau_{\rm DRW}$ on timescales from one hour to one day (see Figure 3), outperforming single-band approaches. When nightly stacked images are required to detect variability from fainter sources, the intra-night cadence in single-band LSST light curves is lost, leading to biased $\tau_{\rm DRW}$ estimates (see Figure 4). In contrast, through multiband fitting, EzTaoX recovers $\tau_{\rm DRW}$ without bias down to hourly timescales for AGNs detected in the WFD survey.
4.5 Caveats & Future Work
Variable emission from extended re-processors is not explicitly modeled in the current implementation of EzTaoX. Recent investigations suggest that reprocessed emission arising at distances much larger than the scale of the accretion disk—for example, in the BLR—can contribute to the measured continuum lags, potentially leading to overestimated accretion disk sizes relative to theoretical predictions (e.g., Korista & Goad, 2001; McHardy et al., 2018; Cackett et al., 2018; Lawther et al., 2018; Lira et al., 2024; Jaiswal et al., 2024; Leighly et al., 2024). The current multiband light curve model in EzTaoX assumes a single lag between any two photometric bands, and therefore lacks the capacity to disentangle multi-component lag distributions, such as reprocessed emission arising from the accretion disk versus that from much larger scales. Developing methodologies that can decompose observed light curves into distinct lagged components would be highly valuable—not only for probing the detailed geometry of AGN accretion flows, but also for advancing our understanding of the physical drivers behind AGN stochastic variability.
The GP kernels currently supported by EzTaoX are all stationary; however, AGN stochastic variability may exhibit non-stationary behavior. A notable example is the class of changing-look AGNs (see Ricci & Trakhtenbrot 2023 and references therein), which exhibit dramatic variability on the timescale of years to decades, likely driven by extreme changes in the accretion rate of the disk (e.g., Ruan et al., 2016; Noda & Done, 2018; Graham et al., 2020; Panda & Śniegowska, 2024). Such systems underscore the limitations of assuming stationarity in modeling AGN variability. Future development of EzTaoX will explore non-stationary GP kernels, in which the kernel parameters evolve with time rather than remaining fixed (e.g., Noack & Sethian, 2021).
5 Summary & Conclusion
We present EzTaoX —a new software package designed to leverage multiband AGN light curves for joint inference of intrinsic stochastic variability properties and interband continuum lags. We evaluate the performance of EzTaoX using both simulated LSST AGN light curves sampled at the cadence of the WFD survey, and multiwavelength light curves of real AGNs. The EzTaoX software is available on GitHub [^7] EzTaoX codebase: https://github.com/ywx649999311/EzTaoX. under a MIT License and version 0.1 is archived in Zenodo (Yu & Burke, 2025). The key advancements and capabilities of EzTaoX are summarized as follows:
  • EzTaoX is $\sim 10^{2}-10^{3}$ times faster than JAVELIN for LSST WFD light curves, and more than $\sim 10^{4}\times$ faster for DDF light curves (see Figure 8 and Section 3.4).
  • EzTaoX and JAVELIN show comparable reliability in recovering simulated interband lags. Specifically, lags can be recovered to within $\sim\,$2 days ($1\sigma$) using the first 3 years of LSST WFD observations (assuming a fiducial $\tau_{\rm DRW}$ of 100 days, a $\sigma_{\mathrm{DRW}}$ of 0.112 mag, and a $g$-band mean magnitude of 21), and to within $\sim\,$0.7 day ($1\sigma$) with the full 10-year dataset (see Figure 2).
  • EzTaoX provides more reliable recovery of the simulated $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ than JAVELIN. In particular, JAVELIN produces systematically biased estimates, whereas EzTaoX recovers unbiased results across the tested range (see Figure 3 and Figure 5).
  • EzTaoX’s multiband modeling capability enables robust recovery of $\tau_{\rm DRW}$ down to hourly timescales with simulated LSST WFD light curves, outperforming single-band modeling. This precision enhances the ability to identify low-mass AGNs (e.g., those in dwarf galaxies) through variability (see Figure 4).
  • EzTaoX supports a broad class of GP kernels beyond the DRW model for joint modeling of AGN stochastic variability and interband lags, offering greater modeling flexibility (see Figure 10 for an example).
In addition to characterizing AGN stochastic variability and measuring continuum lags, EzTaoX is applicable to a variety of other time-domain astrophysical studies. For example, EzTaoX can be applied to measure BLR emission line lags from spectroscopic RM light curves (Blandford & McKee, 1982), or to detect negative lags that may arise from inward propagation of accretion rate fluctuations (Secunda et al., 2023)—both of which typically occur on timescales much longer than the continuum lags. Moreover, EzTaoX can be used to measure time delays between the multiple images of gravitationally lensed quasars, providing an important avenue for constraining the Hubble constant (Blandford & Narayan, 1992). Finally, interband lags have been shown to correlate with AGN continuum luminosity (e.g., Sergeev et al., 2005; Netzer, 2021; Guo et al., 2022; Panda et al., 2024); once properly calibrated, this lag-luminosity relation may enable the estimation of AGN physical properties using photometric time-domain data alone.
We thank Maurizio Paolillo for providing comments to an early draft of this manuscript. W.Y. acknowledges support from the Dunlap Institute for Astronomy & Astrophysics at the University of Toronto. J.J.R. acknowledges support from the Canada Research Chairs (CRC) program, the NSERC Discovery Grant program, the Canada Foundation for Innovation (CFI), and the Québec Ministère de l’Économie et de l’Innovation. C.J.B. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303803. This material is based upon work supported by the National Science Foundation under Award No. 2303803. This research award to NSF is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time. R.J.A., F.E.B, and S.S.S. gratefully acknowledge support from ANID: CATA BASAL FB210003 (R.J.A., F.E.B., S.S.S.); Millennium Science Initiative AIM23-0001 (F.E.B.); and FONDECYT Regular 1231718 (R.J.A.) and 1241005 (F.E.B.). D.D. acknowledges PON R&I 2021, CUP E65F21002880003, and Fondi di Ricerca di Ateneo (FRA), linea C, progetto TORNADO. L.H.G. acknowledges financial support from ANID programs: FONDECYT Iniciación 11241477, Millennium Science Initiative Programs NCN2023-002, and AIM23-0001. D.I. and A.B.K. acknowledge funding provided by the University of Belgrade - Faculty of Mathematics (the contract 451-03-136/2025-03/200104) through the grant by the Ministry of Science, Technological Development and Innovation of the Republic of Serbia. S.P. is supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America.
Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grants No. AST-1440341 and AST-2034437 and a collaboration including current partners Caltech, IPAC, the Oskar Klein Center at Stockholm University, the University of Maryland, University of California, Berkeley , the University of Wisconsin at Milwaukee, University of Warwick, Ruhr University, Cornell University, Northwestern University and Drexel University. Operations are conducted by COO, IPAC, and UW.
Appendix A Additional EzTaoX Experiments
We perform additional experiments to demonstrate the robustness of EzTaoX in recovering DRW parameter under different regimes. To test its performance in recovering $\tau_{\rm DRW}$ at lower intrinsic variability amplitudes, we simulate a set of light curves similar to those in Section 3.1.3, but with the $g$-band $\sigma_{\mathrm{DRW}}$ reduced to 0.06 mag—half the value adopted in Section 3.1.3 and only about four times the average photometric uncertainty. As shown in Figure 12, EzTaoX still robustly recovers $\tau_{\rm DRW}$ down to one hour with 3 years of LSST WFD light curves. To evaluate its performance in recovering $\sigma_{\mathrm{DRW}}$ given larger intrinsic $\tau_{\rm DRW}$, we simulate another set of light curves similar to those in Section 3.1.4, but with the $g$-band $\tau_{\rm DRW}$ increased to 200 days—twice the value adopted in Section 3.1.4. As shown in Figure 13, EzTaoX continues to provide unbiased estimates of $\sigma_{\mathrm{DRW}}$ across the full simulated range.
Figure 12: Recovered versus input $\tau_{\rm DRW}$ from multiband ($u$, $g$, and $r$) light curves with EzTaoX. The 3 year LSST WFD $g$-band light curves are simulated with $\tau_{\rm DRW}$ =100=100 days and $\sigma_{\mathrm{DRW}}$ == 0.06 mag (half the amplitude of the light curves in Figure 3). Even when the intrinsic $\sigma_{\mathrm{DRW}}$ is only a few times the photometric uncertainty, EzTaoX robustly recovers the input $\tau_{\rm DRW}$ down to one hour.
Figure 13: Recovered versus input $\sigma_{\mathrm{DRW}}$ from multiband ($u$, $g$, and $r$) light curves with EzTaoX. The 3 year LSST WFD $g$-band light curves are simulated with $\tau_{\rm DRW}$ =200=200 days and $\sigma_{\mathrm{DRW}}$ == 0.112 mag. Consistent with Figure 5, EzTaoX yields unbiased estimates of $\sigma_{\mathrm{DRW}}$ across the full simulated range.
Appendix B Additional JAVELIN Experiments
The distribution of DRW parameters recovered by JAVELIN’s multiband fitting depends on the properties of the light curves (e.g., cadence and mean photometric uncertainty) used in its initial fitting step, where a prior distribution of $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ is generated from a selected single-band light curve. To illustrate this sensitivity, we refit the light curves from Section 3.1.3 and Section 3.1.4, generating the prior distributions from the $u$- and $g$-band light curves. Figures 14 and 15 show the resulting distributions of $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ recoveries, highlighting that the final DRW parameter distributions differ depending on the band used to generate the priors. Moreover, more than 40% of JAVELIN fits fail the cleaning cuts when the $u$-band is used to generate the priors. This likely arises because the $u$-band has the poorest temporal sampling and highest photometric uncertainty, making the priors from JAVELIN’s first fitting step unreliable.
Figure 14: Recovered versus input $\tau_{\rm DRW}$ from multiband ($u$, $g$, and $r$) light curves with JAVELIN. The blue, green, and red regions show the central 68% intervals for the distributions of the recovered $\tau_{\rm DRW}$ with the priors derived from $u$-, $g$-, and $r$-band light curves, respectively. The recovered $\tau_{\rm DRW}$ distributions differ depending on the band used to generate the priors.
Figure 15: Recovered versus input $\sigma_{\mathrm{DRW}}$ from multiband ($u$, $g$, and $r$) light curves with JAVELIN. The blue, green, and red regions show the central 68% intervals for the distributions of the recovered $\sigma_{\mathrm{DRW}}$ with the priors derived from $u$-, $g$-, and $r$-band light curves, respectively. The recovered $\sigma_{\mathrm{DRW}}$ distributions differ depending on the band used to generate the priors.

### v2

@@
 In addition, EzTaoX’s multiband capability allows robust characterization of AGN stochastic variability down to hourly timescales, facilitating the identification of accreting low-mass AGNs—such as those residing in dwarf galaxies—through their distinctive variability signatures.
-software: pandas (McKinney, 2010), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), astropy (The Astropy Collaboration et al., 2013; The Astropy Collaboration et al., 2018; Astropy Collaboration et al., 2022), emcee (Foreman-Mackey et al., 2013), eztao (Yu & Richards, 2022), tinygp (Foreman-Mackey et al., 2024), numpyro (Bingham et al., 2019; Phan et al., 2019) eztaox (Yu & Burke, 2025)
+software: pandas (pandas2010), numpy (numpy2020), scipy (scipy2020), matplotlib (matplotlib2007), astropy (astropy2013; theastropycollaboration2018; astropycollaboration2022), emcee (foreman-mackey2013), eztao (yu2022), tinygp (foreman-mackey2024), numpyro (bingham2019pyro; phan2019composable) eztaox (eztaox)
 1 Motivation
-The UV/optical continuum luminosity of active galactic nuclei (AGNs) exhibit correlated stochastic variability on timescales ranging from hours to years (e.g., Vanden Berk et al., 2004; Sesar et al., 2007; Schmidt et al., 2012).
-This variability is believed to originate in the accretion disk, potentially driven in part by thermal reprocessing of X-ray photons emitted from the central corona (Krolik et al., 1991) and/or accretion rate fluctuations propagating through the disk (Lyubarskii, 1997).
-For a recent review of the observational properties of AGN variability in the UV/optical and X-ray bands, we refer the reader to Paolillo & Papadakis (2025).
-A damped random walk (DRW) process is commonly adopted to model AGN UV/optical light curves (Kelly et al., 2009).
+The UV/optical continuum luminosity of active galactic nuclei (AGNs) exhibit correlated stochastic variability on timescales ranging from hours to years (e.g., vandenberk2004; sesar2007; schmidt2012).
+This variability is believed to originate in the accretion disk, potentially driven in part by thermal reprocessing of X-ray photons emitted from the central corona (Krolik1991) and/or accretion rate fluctuations propagating through the disk (lyubarskii1997).
+For a recent review of the observational properties of AGN variability in the UV/optical and X-ray bands, we refer the reader to Paolillo2025.
+A damped random walk (DRW) process is commonly adopted to model AGN UV/optical light curves (kelly2009).
 The power spectral density (PSD) of a DRW exhibits a broken power-law (PL; $1/f^{\alpha}$) form, where the PL exponent $\alpha$ transitions from zero at low frequencies to two at high frequencies.
@@
 The asymptotic variability amplitude of the process is denoted by $\sigma_{\mathrm{DRW}}$.
-Although the DRW process is a phenomenological model and the precise PSD shape of intrinsic AGN UV/optical variability remains under debate (e.g., Mushotzky et al., 2011; Kasliwal et al., 2015; Smith et al., 2018; Arévalo et al., 2024; Beard et al., 2025; Petrecca et al., 2024; Yu et al., 2025), it continues to be a valuable and widely used tool for modeling AGN light curves because of its adequacy and computational efficiency.
+Although the DRW process is a phenomenological model and the precise PSD shape of intrinsic AGN UV/optical variability remains under debate (e.g., mushotzky2011; kasliwal2015; smith2018; Arevalo2024; Beard2025; Petrecca2024; Yu2025), it continues to be a valuable and widely used tool for modeling AGN light curves because of its adequacy and computational efficiency.
 Modeling AGN UV/optical light curves as a DRW has revealed compelling correlations between variability and AGN fundamental properties.
-Numerous previous investigations have shown that the best-fit $\sigma_{\mathrm{DRW}}$ correlates negatively with the Eddington ratio ($L/L_{\rm Edd}$) and bolometric luminosity ($L_{\rm bol}$) of AGNs (e.g., MacLeod et al., 2010; Sánchez-Sáez et al., 2018; Suberlak et al., 2021).
-Additionally, the characteristic timescale of the DRW model ($\tau_{\rm DRW}$) correlates positively with $M_{\rm{BH}}$ (e.g., Burke et al., 2021; Wang et al., 2023).
+Numerous previous investigations have shown that the best-fit $\sigma_{\mathrm{DRW}}$ correlates negatively with the Eddington ratio ($L/L_{\rm Edd}$) and bolometric luminosity ($L_{\rm bol}$) of AGNs (e.g., macleod2010; sanchez-saez2018; suberlak2021).
+Additionally, the characteristic timescale of the DRW model ($\tau_{\rm DRW}$) correlates positively with $M_{\rm{BH}}$ (e.g., burke2021; Wang2023a).
 These empirical relations not only provide valuable insights into the physical mechanisms driving AGN variability, but also form the basis for new techniques to estimate AGN fundamental properties using photometric data alone.
-For example, Burke et al.
-(2022) leveraged the $\tau_{\rm DRW}$$M_{\rm{BH}}$ correlation to search for low-mass AGNs powered by accreting massive black holes (MBH; $10^{5}\,$$M_{\odot}$$M_{\rm{BH}}$$10^{7}\,$$M_{\odot}$) at the centers of dwarf galaxies  (i.e., $M_{\star}\lesssim 10^{10}\,M_{\odot}$; Greene et al., 2020).
-AGN UV/optical variability at longer wavelengths is also often observed to lag behind the variability at shorter wavelengths (Collier et al., 1999; Sergeev et al., 2005).
-Under the thermal reprocessing scenario, longer-wavelength photons originate from disk annuli located farther from the central supermassive black hole (SMBH) than those emitting shorter-wavelength photons (Shakura & Sunyaev, 1973).
-Consequently, the reprocessed emission at longer wavelengths exhibits longer light-crossing time delays relative to the central irradiating source (Cackett et al., 2007).
+For example, burke2022 leveraged the $\tau_{\rm DRW}$$M_{\rm{BH}}$ correlation to search for low-mass AGNs powered by accreting massive black holes (MBH; $10^{5}\,$$M_{\odot}$$M_{\rm{BH}}$$10^{7}\,$$M_{\odot}$) at the centers of dwarf galaxies  (i.e., $M_{\star}\lesssim 10^{10}\,M_{\odot}$; greene2020).
+AGN UV/optical variability at longer wavelengths is also often observed to lag behind the variability at shorter wavelengths (Collier1999; sergeev2005).
+Under the thermal reprocessing scenario, longer-wavelength photons originate from disk annuli located farther from the central supermassive black hole (SMBH) than those emitting shorter-wavelength photons (shakura1973).
+Consequently, the reprocessed emission at longer wavelengths exhibits longer light-crossing time delays relative to the central irradiating source (Cackett2007).
 This specific geometry—in which a compact source near the SMBH (e.g., an X-ray corona) irradiates the accretion disk—is commonly referred to as the lamp-post model.
-In this framework, the size of the accretion disk can be inferred by measuring the time lags between UV/optical light curves at different wavelengths—a technique known as continuum reverberation mapping (CRM), which probes the geometry of unresolved sources through time-delay measurements (see Cackett et al., 2021, and references therein).
-Numerous CRM experiments have reported that AGN accretion disk sizes are systematically larger than those predicted by the standard thin-disk model (e.g., Fausnaugh et al., 2016; Jiang et al., 2017; Jha et al., 2022).
-Larger-than-expected accretion disk sizes have also been independently inferred from microlensing observations of gravitationally lensed quasars, where the disk half-light radius (i.e., the characteristic size) is inferred from the differential microlensing variability exhibited by the multiple lensed images (e.g., Pooley et al., 2007; Dai et al., 2010; Morgan et al., 2010; Jiménez-Vicente et al., 2014; Cornachione et al., 2020).
-It has been suggested that the measured lags may be contaminated by reprocessed UV/optical photons originating in the broad emission-line region (BLR), which lies farther out in the accretion flow (e.g., Korista & Goad, 2001; McHardy et al., 2018; Cackett et al., 2018; Lawther et al., 2018).
+In this framework, the size of the accretion disk can be inferred by measuring the time lags between UV/optical light curves at different wavelengths—a technique known as continuum reverberation mapping (CRM), which probes the geometry of unresolved sources through time-delay measurements (see cackett2021, and references therein).
+Numerous CRM experiments have reported that AGN accretion disk sizes are systematically larger than those predicted by the standard thin-disk model (e.g., fausnaugh2016; jiang2017; Jha2022).
+Larger-than-expected accretion disk sizes have also been independently inferred from microlensing observations of gravitationally lensed quasars, where the disk half-light radius (i.e., the characteristic size) is inferred from the differential microlensing variability exhibited by the multiple lensed images (e.g., Pooley2007; Dai2010; morgan2010; Jimenez-Vicente2014; Cornachione2020).
+It has been suggested that the measured lags may be contaminated by reprocessed UV/optical photons originating in the broad emission-line region (BLR), which lies farther out in the accretion flow (e.g., korista2001; mchardy2018; cackett2018; lawther2018).
 This contamination could cause the observed interband continuum lags to appear longer than those expected from pure disk reprocessing.
-In contrast, some studies have argued that accretion disk sizes inferred from CRM are consistent with the standard thin-disk model when factors such as temperature fluctuations within the disk or potential selection biases are properly accounted for (e.g., Dexter & Agol, 2011; Mudd et al., 2018; Homayouni et al., 2019; Yu et al., 2020b).
-Moreover, Kammoun et al.
-(2021) showed that the standard thin-disk model can reproduce the observed CRM lags when relativistic (special and general) and disk ionization effects are fully included.
+In contrast, some studies have argued that accretion disk sizes inferred from CRM are consistent with the standard thin-disk model when factors such as temperature fluctuations within the disk or potential selection biases are properly accounted for (e.g., dexter2011a; mudd2018; homayouni2019; yu2020).
+Moreover, Kammoun2021 showed that the standard thin-disk model can reproduce the observed CRM lags when relativistic (special and general) and disk ionization effects are fully included.
 The accretion disk size discrepancy poses one of the significant challenges to the standard thin-disk model and to our understanding of AGN accretion physics more broadly.
 A large sample of AGNs with robustly measured CRM lags is required to investigate this issue further.
-To date, however, CRM lags have been successfully measured for only a few hundred AGNs (e.g., Collier et al., 1998; Peterson et al., 1998; Sergeev et al., 2005; Jiang et al., 2017; Homayouni et al., 2019; Mudd et al., 2018; Yu et al., 2020b; Guo et al., 2022).
+To date, however, CRM lags have been successfully measured for only a few hundred AGNs (e.g., Collier1998; Peterson1998; sergeev2005; jiang2017; homayouni2019; mudd2018; yu2020; guo2022).
 […]
 The high-quality light curves provided by LSST will enable robust characterization of AGN UV/optical variability and interband continuum lags.
-Interband lags can be measured for about one million LSST AGNs (Yu et al., 2020a, Li et al.
+Interband lags can be measured for about one million LSST AGNs (yu2020a, Li et al.
 2025 submitted), allowing detailed investigations of the accretion disk size problem as a function of AGN properties such as $L_{\rm Edd}$ and $M_{\rm BH}$, and offering new insights into the physical origin of continuum lags.
-Meanwhile, stochastic variability parameters—such as characteristic timescales and amplitudes—extracted from LSST light curves will enable reliable AGN classification (e.g., Kozłowski et al., 2010; Butler & Bloom, 2011; MacLeod et al., 2011; Peters et al., 2015; De Cicco et al., 2021; Sánchez-Sáez et al., 2021; Savić et al., 2023; Sánchez-Sáez et al., 2023).
-For example, variability-based methods are more effective than SED-based methods at identifying low-luminosity and low-mass AGNs (e.g., in dwarf galaxies), where host-galaxy starlight contamination is significant (e.g., Baldassare et al., 2017; Ward et al., 2022; Bernal et al., 2025).
-A complete census of low-luminosity is essential for understanding AGN evolution across cosmic time (Richards et al., 2006), while a comprehensive inventory of low-mass AGNs is critical for constraining the formation pathways of SMBHs (see Greene et al., 2020, and references therein).
+Meanwhile, stochastic variability parameters—such as characteristic timescales and amplitudes—extracted from LSST light curves will enable reliable AGN classification (e.g., kozlowski2010; butler2011; macleod2011; peters2015; DeCicco2021; Sanchez-Saez2021; savic2023; Sanchez-Saez2023).
+For example, variability-based methods are more effective than SED-based methods at identifying low-luminosity and low-mass AGNs (e.g., in dwarf galaxies), where host-galaxy starlight contamination is significant (e.g., baldassare2017; ward2022; bernal2025).
+A complete census of low-luminosity is essential for understanding AGN evolution across cosmic time (richards2006a), while a comprehensive inventory of low-mass AGNs is critical for constraining the formation pathways of SMBHs (see greene2020, and references therein).
 However, existing AGN light curve analysis tools are not well-prepared to fully exploit the scale and richness of LSST observations.
-Current tools/methods used to characterize AGN UV/optical variability typically operate on single-band light curves and lack the ability to incorporate multiband information (e.g., Vaughan et al., 2003; Kelly et al., 2009, 2014; Kozłowski, 2016; Yu & Richards, 2022).
+Current tools/methods used to characterize AGN UV/optical variability typically operate on single-band light curves and lack the ability to incorporate multiband information (e.g., Vaughan2003; kelly2009; kelly2014; kozlowski2016; yu2022).
 This limitation prevents the effective utilization of LSST’s multiband coverage for improved characterization of AGN variability.
 Tools/methods commonly used to measure interband lags generally fall into two categories.
-The first includes techniques that estimate lags by computing the cross-correlation function (CCF) between light curves; the CCF reaches its maximum when one light curve is shifted relative to the other by an amount corresponding to the intrinsic lag (e.g., Gaskell & Peterson, 1987; Edelson & Krolik, 1988; Alexander, 1997).
-[^1] When applied to a single-band light curve and its shifted copies, the CCF method yields its autocorrelation function (ACF), which can be used to characterize AGN variability (e.g., Raiteri et al., 2021a, b).
+The first includes techniques that estimate lags by computing the cross-correlation function (CCF) between light curves; the CCF reaches its maximum when one light curve is shifted relative to the other by an amount corresponding to the intrinsic lag (e.g., gaskell1987; edelson1988; Alexander1997).
+[^1] When applied to a single-band light curve and its shifted copies, the CCF method yields its autocorrelation function (ACF), which can be used to characterize AGN variability (e.g., Raiteri2021a; Raiteri2021b).
 These methods are typically computationally efficient, but are limited to lag estimation and do not provide a full characterization of the underlying variability process.
-The second consists of methods that model each observed light curve as the convolution of an unknown driving continuum with a time-lag distribution, represented by a transfer function (e.g., Zu et al., 2011; Li et al., 2016; Starkey et al., 2016; Donnan et al., 2021).
+The second consists of methods that model each observed light curve as the convolution of an unknown driving continuum with a time-lag distribution, represented by a transfer function (e.g., zu2011; li2016; starkey2016; donnan2021).
 This is expressed as,
@@
 where $y_{\rm drive}(t-\tau)$ denotes the flux of the underlying driving continuum at $t-\tau$, and $\Psi(\tau)$ represents the corresponding transfer function.
-JAVELIN is a representative tool in this category (Zu et al., 2013).
+JAVELIN is a representative tool in this category (zu2013).
 It models the driving continuum as a DRW process and adopts a top-hat profile for the transfer function.
 While such methods provide a more comprehensive treatment by jointly modeling variability and interband lags, they are often computationally demanding.
-Many recent investigations have turned to machine learning (ML) techniques to develop more scalable methods for analyzing AGN light curves (e.g., Fagin et al., 2024; Li et al., 2024).
+Many recent investigations have turned to machine learning (ML) techniques to develop more scalable methods for analyzing AGN light curves (e.g., fagin2024; li2024).
 In these approaches, deep neural networks are trained on simulated light curves and are tasked with recovering the underlying variability characteristics and/or the physical parameters used to generate the simulations.
 However, the reliability of such methods hinges on the assumption that the models used to generate the training data accurately represent both the variability physics and the observational conditions (e.g., cadence, photometric uncertainty)—an assumption rarely satisfied in practice.
-For example, Li et al.
-(2024) show that the AGN physical parameters (e.g., $M_{\rm{BH}}$) predicted by their simulation-based inference model (Cranmer et al., 2020) can be biased by up to 0.5 dex when the test light curves are five times noisier than those used for training.
-In this paper, we present EzTaoX—a scalable and flexible AGN light curve modeling tool that leverages LSST’s multiband observations to simultaneously infer the properties of the underlying stochastic variability and the lags between photometric bands (Yu & Burke, 2025).
+For example, li2024 show that the AGN physical parameters (e.g., $M_{\rm{BH}}$) predicted by their simulation-based inference model (cranmer2020) can be biased by up to 0.5 dex when the test light curves are five times noisier than those used for training.
+In this paper, we present EzTaoX—a scalable and flexible AGN light curve modeling tool that leverages LSST’s multiband observations to simultaneously infer the properties of the underlying stochastic variability and the lags between photometric bands (eztaox).
 The name of EzTaoX combines three elements: “Ez", a homophone of easy; “Tao" referencing the Chinese philosophical and religious tradition of Taoism, commonly translated as “the Way"; and “X" denoting that the code is implemented in JAX—a high-performance numerical computing library developed by Google.
 EzTaoX achieves a speed increase of $\sim 10^{2}-10^{4}\times$ on CPUs compared to existing tools with similar capabilities (e.g., JAVELIN), making it feasible to perform both stochastic variability characterization and interband continuum lag measurements for all AGNs discovered by LSST.
-Furthermore, EzTaoX’s multiband capability enables robust constraints on stochastic variability timescales (e.g., $\tau_{\rm DRW}$) down to one hour with the LSST WFD cadence, facilitating the identification of low-mass AGNs—such as those residing in dwarf galaxies—which are expected to exhibit characteristic $\tau_{\rm DRW}$ ranging from hours to days (Burke et al., 2021; Wang et al., 2023).
+Furthermore, EzTaoX’s multiband capability enables robust constraints on stochastic variability timescales (e.g., $\tau_{\rm DRW}$) down to one hour with the LSST WFD cadence, facilitating the identification of low-mass AGNs—such as those residing in dwarf galaxies—which are expected to exhibit characteristic $\tau_{\rm DRW}$ ranging from hours to days (burke2021; Wang2023a).
 The rest of the paper is organized as follows.
@@
 2 Light Curve Modeling Method & Implementation
-EzTaoX models AGN UV/optical stochastic variability as a Gaussian processes (GP), a method commonly adopted in AGN variability analysis (e.g., Press et al., 1992; Kelly et al., 2009; Kozłowski et al., 2010; Wilkins, 2019; Griffiths et al., 2021; Yu et al., 2022; Stone et al., 2022; Zhang et al., 2023; McLaughlin et al., 2024).
-Since the flux light curves of accreting compact objects typically follow a lognormal distribution (e.g., Uttley & McHardy, 2001; Gaskell, 2004; Giebels & Degrange, 2009; MacLeod et al., 2012; De Cicco et al., 2022), we emphasize that GP modeling is best applied to magnitude light curves, which follow a Gaussian distribution.
-[^2] Gúrpide & Middleton (2025) demonstrated that even when the light curves follow a lognormal distribution, the simulated GP parameters can be recovered without significant bias.
-Given a single band AGN light curve, a GP treats the intrinsic brightness of the AGN at each point in time as a Gaussian random variable and captures the temporal correlations between observations through a kernel function $k(t_{i},t_{j})$, which defines the covariance between the Gaussian variables at any two timestamps $t_{i}$ and $t_{j}$ (Rasmussen & Williams, 2006; Aigrain & Foreman-Mackey, 2023).
+EzTaoX models AGN UV/optical stochastic variability as a Gaussian process (GP), a method commonly adopted in AGN variability analysis (e.g., Press1992; kelly2009; kozlowski2010; Wilkins2019; Griffiths2021; yu2022b; stone2022; Zhang2023; McLaughlin2024).
+Since the flux light curves of accreting compact objects typically follow a lognormal distribution (e.g., uttley2001; gaskell2004; Giebels2009; macleod2012; decicco2022), we emphasize that GP modeling is best applied to magnitude light curves, which follow a Gaussian distribution.
+Moreover, Gurpide2025 demonstrated that even when the light curves follow a lognormal distribution, the simulated GP parameters can be recovered without significant bias.
+Given a single band AGN light curve, a GP treats the intrinsic brightness of the AGN at each point in time as a Gaussian random variable and captures the temporal correlations between observations through a kernel function $k(t_{i},t_{j})$, which defines the covariance between the Gaussian variables at any two timestamps $t_{i}$ and $t_{j}$ (Rasmussen2006; aigrain2023).
 The likelihood function of a GP given a set of observations $\bf{y}$ is,
@@
 The variability amplitude in each band is obtained by multiplying the latent GP amplitude by the corresponding scaling factor.
-This class of kernel functions, as expressed in Equation 5, has seen broad application in astronomical time series analysis (e.g., Gordon et al., 2020; Villar et al., 2021).
+This class of kernel functions, as expressed in Equation 5, has seen broad application in astronomical time series analysis (e.g., gordon2020; Villar2021).
 To evaluate the likelihood of the multiband model for a given light curve, we shift the light curve in different bands by their trial $\tau_{\rm lag}$ relative to the reference band, and compute Equation 2 using a kernel that assumes no intrinsic lag between bands (i.e., setting $\tau_{\rm lag}=0$ in Equation 5).
@@
 [^3] https://github.com/jax-ml/jax JAX is a high-performance numerical computing library that combines automatic differentiation, just-in-time (JIT) compilation, and hardware acceleration to empower scalable machine learning and scientific computing.
-tinygp leverages these features of JAX and re-implements the novel celerite algorithm introduced by Foreman-Mackey et al.
-(2017) in JAX.
-When a kernel function can be expressed as a mixture of exponentials, the celerite algorithm can evaluate the GP likelihood function with $\mathcal{O}(N)$ computational scaling, a significant efficiency enhancement over the standard $\mathcal{O}(N^{3})$ scaling associated with direct likelihood evaluation (Foreman-Mackey et al., 2017).
+tinygp leverages these features of JAX and re-implements the novel celerite algorithm introduced by foreman-mackey2017 in JAX.
+When a kernel function can be expressed as a mixture of exponentials, the celerite algorithm can evaluate the GP likelihood function with $\mathcal{O}(N)$ computational scaling, a significant efficiency enhancement over the standard $\mathcal{O}(N^{3})$ scaling associated with direct likelihood evaluation (foreman-mackey2017).
 EzTaoX takes advantage of the JAX-based implementation of the celerite algorithm to provide a scalable solution for multiband light curve modeling of AGNs.
 PSD shapes exhibited by the light curves of accreting black holes can be effectively modeled with celerite GP kernels.
-For example, the continuous-time autoregressive moving-average (CARMA) process—commonly used as a more flexible alternative to the DRW process—has an equivalent celerite implementation (Roux, 2002; Kelly et al., 2014; Kasliwal et al., 2017; Moreno et al., 2019; Yu et al., 2022).
+For example, the continuous-time autoregressive moving-average (CARMA) process—commonly used as a more flexible alternative to the DRW process—has an equivalent celerite implementation (rouxalet2002; kelly2014; kasliwal2017; moreno2019; yu2022b).
 A CARMA($p$, $q$) process combines an autoregressive process of order $p$ with a moving-average process of order $q$, and its PSD can be expressed as a weighted sum of modified Lorentzians.
 The DRW process is a special case of a CARMA process with $p=1$ and $q=0$.
-For a detailed description of CARMA modeling, we refer the reader to Kelly et al.
-(2014).
-More recently, Lefkir et al.
-(2025) introduced an algorithm for approximating arbitrary broken PL PSDs using a sum of multiple celerite kernels, thereby enhancing the interpretability and flexibility of GP-based modeling for AGN light curves.
+For a detailed description of CARMA modeling, we refer the reader to kelly2014.
+More recently, Lefkir2025 introduced an algorithm for approximating arbitrary broken PL PSDs using a sum of multiple celerite kernels, thereby enhancing the interpretability and flexibility of GP-based modeling for AGN light curves.
 Moreover, celerite-based kernels can be freely combined through addition and multiplication to model more complex covariance structures, while retaining the $\mathcal{O}(N)$ computational scaling.
@@
 We adopt the DRW model to simulate AGN light curves in individual LSST bands.
-We start by simulating DRW light curves in each LSST band with a 30-second cadence; this 30-second cadence is chosen to match the 30-second exposure time of each LSST single visit (Ivezić et al., 2019).
-The input DRW parameters across different bands are correlated according to the wavelength-dependent power-law relations presented in MacLeod et al.
-(2010).
+We start by simulating DRW light curves in each LSST band with a 30-second cadence; this 30-second cadence is chosen to match the 30-second exposure time of each LSST single visit (ivezic2019).
+The input DRW parameters across different bands are correlated according to the wavelength-dependent power-law relations presented in macleod2010.
 More specifically,
 $\tau_{\rm DRW}\propto\lambda^{0.17},\quad\sigma_{\rm DRW}\propto\lambda^{-0.479},$ (6)
-where $\lambda$ is the effective wavelength of the LSST filters (Ivezić et al., 2019).
+where $\lambda$ is the effective wavelength of the LSST filters (ivezic2019).
 We then simulate continuum lags by offsetting light curves in redder band in time with respect to light curves in bluer bands.
@@
 where $\sigma_{\mathrm{sys}}$ is the systematic photometric error and $\sigma_{\mathrm{rand}}$ is the random photometric error.
-LSST aims to achieve a $\sigma_{\mathrm{sys}}<0.005$ mag (Ivezić et al., 2019); we set $\sigma_{\mathrm{sys}}$ to $0.005$ mag in our simulation.
+LSST aims to achieve a $\sigma_{\mathrm{sys}}<0.005$ mag (ivezic2019); we set $\sigma_{\mathrm{sys}}$ to $0.005$ mag in our simulation.
 The random photometric error ($\sigma_{\mathrm{rand}}$) is expected to scale as
@@
 […]
-The observing cadence and the expected $m_{5}$ at any location on the sky are predicted by LSST operation simulations (OpSims) using MAF (Marshall et al., 2017).
+The observing cadence and the expected $m_{5}$ at any location on the sky are predicted by LSST operation simulations (OpSims) using MAF (marshall2017).
 In this work, we adopt the operation simulation version baseline_v4.0_10yrs.
-All our simulated AGNs also assume a fiducial $g$-band mean magnitude of 21 ($\sim 1\%$ photometric accuracy) and mean uugg and ggrr colors of 0.1 mag (Richards et al., 2001; Temple et al., 2021).
+All our simulated AGNs also assume a fiducial $g$-band mean magnitude of 21 ($\sim 1\%$ photometric accuracy) and mean uugg and ggrr colors of 0.1 mag (richards2001; temple2021).
 An example light curve from the first set is shown in Figure 1.
@@
 The ggrr lag for each multiband light curve is uniformly sampled between 0.01 day and 10 days, while the input $g$-band $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ are fixed at fiducial values of 100 days and 0.112 mag, respectively.
-These choices of $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ are representative of those measured from typical quasar (i.e., high luminosity AGN) UV/optical light curves (e.g., Vanden Berk et al., 2004; MacLeod et al., 2010).
+These choices of $\tau_{\rm DRW}$ and $\sigma_{\mathrm{DRW}}$ are representative of those measured from typical quasar (i.e., high luminosity AGN) UV/optical light curves (e.g., vandenberk2004; macleod2010).
 The second set tests EzTaoX’s ability to recover the DRW characteristic timescale $\tau_{\rm DRW}$ (see Section 3.1.3) , and the light curves consist of simulated LSST observations in $u$, $g$, and $r$ bands.
@@
 For example, the lower limit for $\tau_{\rm DRW}$ is set to $\sim$10 times smaller than the minimum time separation between any two observations, while the upper limit is $\sim$10 times larger than the light curve length.
-MCMC sampling is performed using the gradient-based No-U-Turn Sampler (NUTS; Hoffman & Gelman, 2011), with a single chain run for $10^{4}$ steps following $10^{3}$ warm-up steps.
+MCMC sampling is performed using the gradient-based No-U-Turn Sampler (NUTS; Hoffman2011), with a single chain run for $10^{4}$ steps following $10^{3}$ warm-up steps.
 The posterior median of $\tau_{\rm DRW}$ is taken as the point estimate.
@@
 In the first step, the light curve in the bluest band is typically used, as JAVELIN assumes that redder-band light curves are convolved versions of it; here we use the $g$-band light curve.
-For MCMC sampling, JAVELIN employs the affine-invariant ensemble sampler (Goodman & Weare, 2010) as implemented in emcee (Foreman-Mackey et al., 2013).
+For MCMC sampling, JAVELIN employs the affine-invariant ensemble sampler (goodman2010) as implemented in emcee (foreman-mackey2013).
 In the first step we follow the default configuration, while in the second step we run the sampler with 100 walkers, 500 burn-in steps, and 3000 sampling steps.
@@
 We clean the sample of best-fit EzTaoX and JAVELIN lags by removing those associated with non-convergent MCMC chains.
-Specifically, we discard fits in which the MCMC chain for $\tau_{\rm gr}$ yields fewer than 400 effective samples ($n_{\rm eff}$), following the recommendation of Vehtari et al.
-(2021) for selecting converged chains.
+Specifically, we discard fits in which the MCMC chain for $\tau_{\rm gr}$ yields fewer than 400 effective samples ($n_{\rm eff}$), following the recommendation of Vehtari2021 for selecting converged chains.
 Approximately 20% of the 3-year and 7% of the 10-year EzTaoX fits are removed by this criterion, while $\sim$20% of the JAVELIN fits are removed from both the 3-year and 10-year data.
@@
 To better demonstrate the advantages of multiband fitting, we refit the second set of simulated light curves after removing revisits occurring within 1 hr in any given band.
-This removal mimics a scenario in which intra-night observations must be stacked to detect faint variable sources against host-galaxy starlight—such as low-luminosity and low-mass AGNs, which are expected to exhibit shorter characteristic variability timescales than typical quasars (e.g., Burke et al., 2021).
+This removal mimics a scenario in which intra-night observations must be stacked to detect faint variable sources against host-galaxy starlight—such as low-luminosity and low-mass AGNs, which are expected to exhibit shorter characteristic variability timescales than typical quasars (e.g., burke2021).
 Figure 4 displays the recovered $\tau_{\rm DRW}$ as a function of the input value.
@@
 The light curves of NGC 5548 are obtained by the AGN STORM project.
-In the legend, F16 refers to the lags reported by Fausnaugh et al.
-(2016), and E15 refers to the lags reported by Edelson et al.
-(2015).
+In the legend, F16 refers to the lags reported by fausnaugh2016, and E15 refers to the lags reported by Edelson2015.
 Interband continuum lags measured with EzTaoX are broadly consistent with those obtained using JAVELIN and the ICCF method, highlighting the effectiveness of EzTaoX despite its adoption of a simplified transfer function (i.e., a Dirac delta function).
-Figure 7: Comparison of interband continuum lags for a sample of type 1 AGNs measured with EzTaoX to those reported in the literature (Guo et al., 2022).
+Figure 7: Comparison of interband continuum lags for a sample of type 1 AGNs measured with EzTaoX to those reported in the literature (guo2022).
 The light curves are obtained from ZTF.
@@
 3.2 A Test Using Continuum Reverberation Mapping Light Curves of NGC 5548
-The AGN Space Telescope and Optical Reverberation Mapping (STORM) project is an intensive multiwavelength monitoring campaign targeting NGC 5548, a well-studied Seyfert galaxy (De Rosa et al., 2015; Edelson et al., 2015; Fausnaugh et al., 2016).
-This campaign combines observations from the Hubble Space Telescope (HST), the Neil Gehrels Swift Observatory (Swift hereafter; Gehrels et al., 2004), and a network of ground-based telescopes, to provide simultaneous daily light curves of NGC 5548 across the X-ray, UV, and optical bands.
-For further details on the observational setup and data, we refer interested readers to De Rosa et al.
-(2015), Edelson et al.
-(2015), and Fausnaugh et al.
-(2016).
+The AGN Space Telescope and Optical Reverberation Mapping (STORM) project is an intensive multiwavelength monitoring campaign targeting NGC 5548, a well-studied Seyfert galaxy (DeRosa2015; Edelson2015; fausnaugh2016).
+This campaign combines observations from the Hubble Space Telescope (HST), the Neil Gehrels Swift Observatory (Swift hereafter; Gehrels2004), and a network of ground-based telescopes, to provide simultaneous daily light curves of NGC 5548 across the X-ray, UV, and optical bands.
+For further details on the observational setup and data, we refer interested readers to DeRosa2015, Edelson2015, and fausnaugh2016.
 Applying EzTaoX to the UV/optical AGN STORM light curves of NGC 5548 reveals that the lags inferred by EzTaoX are broadly consistent with literature values derived using JAVELIN and ICCF.
 We follow the same strategy described in Section 3.1.2 to measure interband lags relative to the continuum light curve at 1367Å.
-In contrast to the approaches taken by Edelson et al.
-(2015), and Fausnaugh et al.
-(2016), we do not de-trend the light curves prior to lag estimation.
+In contrast to the approaches taken by Edelson2015, and fausnaugh2016, we do not de-trend the light curves prior to lag estimation.
 We also increased the prior range of $S_{\rm band}$ to Uniform($^{-5}, 1010$) because AGN STORM light curve are provided in flux, which exhibits large amplitude variation across bands.
 Figure 6 compares the lags measured with EzTaoX to those derived with JAVELIN and ICCF.
-F16 and E15 in the legend refer to the lags reported by Fausnaugh et al.
-(2016) and Edelson et al.
-(2015), respectively.
+F16 and E15 in the legend refer to the lags reported by fausnaugh2016 and Edelson2015, respectively.
 The blue points represent the continuum lags measured at 1158Å, 1479Å, and 1746Å  using HST light curves.
 The green points correspond to the continuum lags from the Swift UVW2, UVM2, UVW1, U, and B bands.
-The yellow points represent continuum lags from Johnson/Cousins BVRI bands and Sloan Digital Sky Survey (SDSS; York et al., 2000) ugrizugriz bands based on ground-based observations.
-The purple squares in the right panel of Figure 6 shows the Swift lags reported by Edelson et al.
-(2015).
-We note that the Swift ICCF lags from Fausnaugh et al.
-(2016) are systematically larger than those reported by Edelson et al.
-(2015), likely due to the different de-trending approaches they adopted.
+The yellow points represent continuum lags from Johnson/Cousins BVRI bands and Sloan Digital Sky Survey (SDSS; york2000) ugrizugriz bands based on ground-based observations.
+The purple squares in the right panel of Figure 6 shows the Swift lags reported by Edelson2015.
+We note that the Swift ICCF lags from fausnaugh2016 are systematically larger than those reported by Edelson2015, likely due to the different de-trending approaches they adopted.
 Overall, interband lags measured using EzTaoX are broadly consistent with those obtained using JAVELIN and ICCF.
@@
 […]
-We apply EzTaoX to the $g$- and $r$-band ZTF light curves of the core-sample AGNs from Guo et al.
-(2022), and find that the lags measured with EzTaoX are broadly consistent with those obtained using JAVELIN and ICCF.
+We apply EzTaoX to the $g$- and $r$-band ZTF light curves of the core-sample AGNs from guo2022, and find that the lags measured with EzTaoX are broadly consistent with those obtained using JAVELIN and ICCF.
 Figure 7 compares the interband lags measured with EzTaoX to those from JAVELIN (left panel) and ICCF (right panel).
-We retrieve the ZTF light curves for these sources from the IPAC Infrared Science Archive (ZTF Team, 2025) and truncate them to match the DR7 temporal baseline (i.e., MJD < 59400).
+We retrieve the ZTF light curves for these sources from the IPAC Infrared Science Archive (ztf_lightcurve) and truncate them to match the DR7 temporal baseline (i.e., MJD < 59400).
 When multiple light curves are available for the same object (each obtained from a different CCD), we select the one with the largest number of epochs.
 We follow the same strategy described in Section 3.1.2 to measure the lags between the $g$ and $r$ bands.
-We also fit additional jitter terms—one for each band—added to the diagonal elements of the covariance matrix to account for any residual photometric uncertainties not captured by the photometric pipeline (Burke et al., 2021).
+We also fit additional jitter terms—one for each band—added to the diagonal elements of the covariance matrix to account for any residual photometric uncertainties not captured by the photometric pipeline (burke2021).
 3.4 Scalability: Computation Time
@@
 Unbiased recovery of $\tau_{\rm DRW}$ depends on the light-curve cadence, baseline (i.e., the total duration of observations), and adopted priors.
-As shown by Kozłowski (2017) and later studies (e.g., Sánchez-Sáez et al., 2018; Burke et al., 2021; Suberlak et al., 2021; Hu et al., 2024), $\tau_{\rm DRW}$ is systematically underestimated when the light-curve baseline is shorter than $\sim$10 times the true $\tau_{\rm DRW}$.
-Hu et al.
-(2024) conducted extensive simulations demonstrating that the assumed prior on $\tau_{\rm DRW}$ also impact its recovery.
+As shown by kozlowski2017a and later studies (e.g., sanchez-saez2018; burke2021; suberlak2021; Hu2024), $\tau_{\rm DRW}$ is systematically underestimated when the light-curve baseline is shorter than $\sim$10 times the true $\tau_{\rm DRW}$.
+Hu2024 conducted extensive simulations demonstrating that the assumed prior on $\tau_{\rm DRW}$ also impact its recovery.
 In addition, they reported that $\tau_{\rm DRW}$ is systematically overestimated when the true $\tau_{\rm DRW}$ is less than five times the mean cadence, since the temporal sampling then probes only the white-noise regime (i.e., a PL exponent of zero) of the DRW PSD.
@@
 As an illustration, Figure 10 presents $g$- and $r$-band light curves simulated using a CARMA(2,1) process, sampled according to the LSST WFD cadence (left), along with the corresponding PSDs recovered using EzTaoX (right).
-The CARMA(2,1) process—also known as the noise-driven damped harmonic oscillator (DHO) model—has been shown to provide superior fits to observed AGN UV/optical light curves compared to the DRW model (e.g., Kasliwal et al., 2017; Yu et al., 2025).
+The CARMA(2,1) process—also known as the noise-driven damped harmonic oscillator (DHO) model—has been shown to provide superior fits to observed AGN UV/optical light curves compared to the DRW model (e.g., kasliwal2017; Yu2025).
 For reference, the DRW PSD is shown in blue in the right panel.
@@
 4.3 Beyond the Dirac Delta Transfer Function
-EzTaoX models interband lags in multiband light curves using a Dirac delta transfer function, a simplified approach compared to methods that adopt more flexible transfer functions with finite width and skewed shapes (e.g., Zu et al., 2011; Starkey et al., 2016; Li et al., 2016; Donnan et al., 2021).
-While more sophisticated transfer functions can provide additional insights into the geometry and physical properties of AGN accretion disks (e.g., Cackett et al., 2007; Starkey et al., 2016; Fagin et al., 2024), we argue that the cadence of LSST WFD light curves is likely insufficient to resolve the detailed shape of the transfer functions.
+EzTaoX models interband lags in multiband light curves using a Dirac delta transfer function, a simplified approach compared to methods that adopt more flexible transfer functions with finite width and skewed shapes (e.g., zu2011; starkey2016; li2016; donnan2021).
+While more sophisticated transfer functions can provide additional insights into the geometry and physical properties of AGN accretion disks (e.g., Cackett2007; starkey2016; fagin2024), we argue that the cadence of LSST WFD light curves is likely insufficient to resolve the detailed shape of the transfer functions.
 Thus, the simplified Dirac delta transfer function adopted by EzTaoX is likely adequate for lag estimation from LSST WFD light curves.
-Furthermore, previous investigations and the results of Section 3.2 show that even with high-cadence continuum reverberation mapping data, the use of more flexible transfer functions does not necessarily result in significantly different mean lag estimates compared to simpler methods (e.g., Fausnaugh et al., 2018; Donnan et al., 2023).
-An exception perhaps arises when the intrinsic transfer function is notably skewed, in which case simplified methods may underestimate the mean lag (Yu et al., 2020a; Chan et al., 2020).
+Furthermore, previous investigations and the results of Section 3.2 show that even with high-cadence continuum reverberation mapping data, the use of more flexible transfer functions does not necessarily result in significantly different mean lag estimates compared to simpler methods (e.g., Fausnaugh2018; donnan2023).
+An exception perhaps arises when the intrinsic transfer function is notably skewed, in which case simplified methods may underestimate the mean lag (yu2020a; Chan2020).
 As an extension to the comparison in Section 3.1.2, we test EzTaoX and JAVELIN on new sets of 10-year LSST WFD light curves simulated with a Gaussian transfer function.
@@
 EzTaoX enables reliable and simultaneous characterization of variability and interband lags for the tens of millions of AGNs to be monitored by LSST.
-Existing non-ML methods/tools capable of jointly characterizing stochastic variability and measuring interband lags are not scalable to LSST data volumes (e.g., Zu et al., 2011; Starkey et al., 2016; Donnan et al., 2021).
+Existing non-ML methods/tools capable of jointly characterizing stochastic variability and measuring interband lags are not scalable to LSST data volumes (e.g., zu2011; starkey2016; donnan2021).
 While ML approaches are scalable, their reliability critically depends on the assumption that the training data accurately capture the true nature of AGN variability—an assumption that is rarely satisfied in practice.
@@
 EzTaoX enables robust variability selection of actively accreting MBHs.
-Accreting MBHs in dwarf galaxies are difficult to identify due to their low luminosity relative to their host galaxy starlight (Greene et al., 2020).
-Nevertheless, their characteristic stochastic variability provides a powerful selection tool in the time domain (e.g., Baldassare et al., 2018; Ward et al., 2022; Burke et al., 2022; Bernal et al., 2025).
-By modeling dwarf galaxy image-differencing light curves as a DRW process, one can identify MBH candidates, which are expected to exhibit characteristic $\tau_{\rm DRW}$ values ranging from hours to days (Burke et al., 2021; Wang et al., 2023).
+Accreting MBHs in dwarf galaxies are difficult to identify due to their low luminosity relative to their host galaxy starlight (greene2020).
+Nevertheless, their characteristic stochastic variability provides a powerful selection tool in the time domain (e.g., baldassare2018; ward2022; burke2022; bernal2025).
+By modeling dwarf galaxy image-differencing light curves as a DRW process, one can identify MBH candidates, which are expected to exhibit characteristic $\tau_{\rm DRW}$ values ranging from hours to days (burke2021; Wang2023a).
 Given the cadence of LSST WFD light curves, the multiband modeling capability of EzTaoX enables precise measurement of $\tau_{\rm DRW}$ on timescales from one hour to one day (see Figure 3), outperforming single-band approaches.
@@
 Variable emission from extended re-processors is not explicitly modeled in the current implementation of EzTaoX.
-Recent investigations suggest that reprocessed emission arising at distances much larger than the scale of the accretion disk—for example, in the BLR—can contribute to the measured continuum lags, potentially leading to overestimated accretion disk sizes relative to theoretical predictions (e.g., Korista & Goad, 2001; McHardy et al., 2018; Cackett et al., 2018; Lawther et al., 2018; Lira et al., 2024; Jaiswal et al., 2024; Leighly et al., 2024).
+Recent investigations suggest that reprocessed emission arising at distances much larger than the scale of the accretion disk—for example, in the BLR—can contribute to the measured continuum lags, potentially leading to overestimated accretion disk sizes relative to theoretical predictions (e.g., korista2001; mchardy2018; cackett2018; lawther2018; Lira2024; Jaiswal2024; Leighly2024).
 The current multiband light curve model in EzTaoX assumes a single lag between any two photometric bands, and therefore lacks the capacity to disentangle multi-component lag distributions, such as reprocessed emission arising from the accretion disk versus that from much larger scales.
@@
 The GP kernels currently supported by EzTaoX are all stationary; however, AGN stochastic variability may exhibit non-stationary behavior.
-A notable example is the class of changing-look AGNs (see Ricci & Trakhtenbrot 2023 and references therein), which exhibit dramatic variability on the timescale of years to decades, likely driven by extreme changes in the accretion rate of the disk (e.g., Ruan et al., 2016; Noda & Done, 2018; Graham et al., 2020; Panda & Śniegowska, 2024).
+A notable example is the class of changing-look AGNs (see Ricci2023 and references therein), which exhibit dramatic variability on the timescale of years to decades, likely driven by extreme changes in the accretion rate of the disk (e.g., Ruan2016; Noda2018; graham2020; Panda2024b).
 Such systems underscore the limitations of assuming stationarity in modeling AGN variability.
-Future development of EzTaoX will explore non-stationary GP kernels, in which the kernel parameters evolve with time rather than remaining fixed (e.g., Noack & Sethian, 2021).
+Future development of EzTaoX will explore non-stationary GP kernels, in which the kernel parameters evolve with time rather than remaining fixed (e.g., Noack2021).
 5 Summary & Conclusion
@@
 The EzTaoX software is available on GitHub [^6] EzTaoX codebase: https://github.com/ywx649999311/EzTaoX.
-under a MIT License and version 0.1 is archived in Zenodo (Yu & Burke, 2025).
+under a MIT License and version 0.1 is archived in Zenodo (eztaox).
 The key advancements and capabilities of EzTaoX are summarized as follows:
@@
 In addition to characterizing AGN stochastic variability and measuring continuum lags, EzTaoX is applicable to a variety of other time-domain astrophysical studies.
-For example, EzTaoX can be applied to measure BLR emission line lags from spectroscopic RM light curves (Blandford & McKee, 1982), or to detect negative lags that may arise from inward propagation of accretion rate fluctuations (Secunda et al., 2023)—both of which typically occur on timescales much longer than the continuum lags.
-Moreover, EzTaoX can be used to measure time delays between the multiple images of gravitationally lensed quasars, providing an important avenue for constraining the Hubble constant (Blandford & Narayan, 1992).
-Finally, interband lags have been shown to correlate with AGN continuum luminosity (e.g., Sergeev et al., 2005; Netzer, 2021; Guo et al., 2022; Panda et al., 2024); once properly calibrated, this lag-luminosity relation may enable the estimation of AGN physical properties using photometric time-domain data alone.
-We thank Maurizio Paolillo for providing comments to an early draft of this manuscript.
+For example, EzTaoX can be applied to measure BLR emission line lags from spectroscopic RM light curves (blandford1982), or to detect negative lags that may arise from inward propagation of accretion rate fluctuations (Secunda2023)—both of which typically occur on timescales much longer than the continuum lags.
+Moreover, EzTaoX can be used to measure time delays between the multiple images of gravitationally lensed quasars, providing an important avenue for constraining the Hubble constant (Blandford1992).
+Finally, interband lags have been shown to correlate with AGN continuum luminosity (e.g., sergeev2005; netzer2021; guo2022; Panda2024a); once properly calibrated, this lag-luminosity relation may enable the estimation of AGN physical properties using photometric time-domain data alone.
+We thank the anonymous referee for the thoughtful comments that helped us improve the paper.
+We thank Maurizio Paolillo and Thaisa Storchi Bergmann for providing comments to an early draft of this manuscript.
 W.Y.